### v1

# Worst-Case to Expander-Case Reductions: Derandomized and Generalized

Abstract
A recent paper by Abboud and Wallheimer [ITCS 2023] presents self-reductions for various fundamental graph problems, that transform worst-case instances to expanders, thus proving that the complexity remains unchanged if the input is assumed to be an expander. An interesting corollary of their self-reductions is that, if some problem admit such reduction, then the popular algorithmic paradigm based on expander-decompositions is useless against it. In this paper, we improve their core gadget, which augments a graph to make it an expander while retaining its important structure. Our new core construction has the benefit of being simple to analyze and generalize, while obtaining the following results:
  • A derandomization of the self-reductions, showing that the equivalence between worst-case and expander-case holds even for deterministic algorithms, and ruling out the use of expander-decompositions as a derandomization tool.
  • An extension of the results to other models of computation, such as the Fully Dynamic model and the Congested Clique model. In the former, we either improve or provide an alternative approach to some recent hardness results for dynamic expander graphs, by Henzinger, Paz, and Sricharan [ESA 2022].
In addition, we continue this line of research by designing new self-reductions for more problems, such as Max-Cut and dynamic Densest Subgraph, and demonstrating that the core gadget can be utilized to lift lower bounds based on the OMv Conjecture to expanders.
1 Introduction
When studying the complexity of any graph problem, it is natural to ask whether the problem can be solved faster on expanders, i.e., random-like, well-connected graphs that satisfy a certain definition of expansion.
Question 1.
Are expanders worst-case instances of my problem?
The motivation for such a question comes from multiple sources. First, it is inherently interesting to understand how the rich mathematical structure of expanders affects the complexity of fundamental problems such as shortest paths, cuts, matchings, subgraph detection, and so on. After all, expanders are among the most important graph families in computer science. Second, expanders exhibit some of the most algorithmically useful properties of uniformly random graphs, and so this question may help understand the average-case complexity. Third, graphs that arise in applications may be expanders (e.g., in network architecture). Moreover last but not least is the hope that if we solve a problem faster on expanders, we will also be able to solve it in the worst case by utilizing the popular expander decomposition method, which we discuss soon. Unless explicitly stated otherwise, we use the conductance-based notion of $\phi$-expanders, whose precise definition can be found in Section 3, and we say that a graph is an expander if it is an $\Omega(1)$-expander.
It is possible to cook up problems for which the answer to Question 1 is negative. For instance, we can solve connectivity in constant time if the input is promised to be an expander, but it requires linear time in the worst case. Less obvious examples with such gaps are counting spanning trees [32] and computing short paths [8] [^2] Using the stronger notion of spectral expanders, which we do not consider here.. However, for many (perhaps most) interesting graph problems, the answer seems to be positive: expanders do not make the problem any easier. In other words, the expander-case is also worst-case. But how do we prove that? Let us discuss three methods and their drawbacks.
  1. 1.
    The first and most obvious method is to prove a lower bound for the problem on expander instances that matches the worst-case upper bound. Technically, this may follow directly from the existing lower bounds for the problem since they are often proved on random-like graphs (e.g., for distance oracles [40]), or it may require some modifications to the lower bound proofs (e.g., for dynamic graph problems [26]).
    The main drawback with this approach is that we are interested in answering Question 1 even when (or rather, especially when) we have not already resolved the worst-case time complexity of our problem, in which case we do not even have a matching lower bound (e.g., the Maximum Matching problem). Another drawback is that when asking for the fine-grained complexity of problems, the existing lower bounds are usually conditioned on strong assumptions, and one may hope to get an unconditional answer to Question 1.
    The next two approaches resolve these drawbacks since they are based on worst-case to expander-case self-reductions (WTERs). Such techniques show equivalence between the expander-case complexity and the worst-case complexity.
  2. 2.
    The second approach uses the expander decomposition method. This is a popular paradigm in recent years that suggests we can solve graph problems by (1) decomposing the graph into vertex-disjoint expanders with a small number of edges between them, (2) solving the problem on each expander separately, and (3) combining all the answers efficiently. Step (3) requires problem-specific techniques. If we can solve steps (1) and (3) for a problem, then we have effectively shown that any improvement on expanders will yield an improvement on worst-case graphs, giving a positive answer to Question 1. Efficient algorithms for computing such expander decompositions (for step (1)) are known both in the Word-RAM  [5, 38, 28, 41, 34] and in other models of computation including dynamic [38], distributed [13], and recently even in streaming [17]. Applications of this paradigm have led to many breakthroughs in recent years to problems such as Maximum Flow [14], Dynamic Connectivity [23], Gomory-Hu Trees [2], Minimum Spanning Trees [33], and Triangle Enumeration [13].
    Self-reductions of this form are called ED-WTERs. A recent paper by Abboud and Wallheimer [3] proposed an alternative, simpler method of self-reductions that do not use expander decompositions. They call this method Direct-WTERs, which we discuss next, and it is not only simpler but also yields stronger qualitative and quantitative results.
  3. 3.
    The third and most direct method is to show that any graph can be turned into an expander without affecting the solution to the problem or increasing the size of the graph by too much. In their paper, Abboud and Wallheimer [3] gave the following definition, which we slightly reframe to fit our discussion more accurately:
    Definition 1 (Direct-WTER [3, Definition 2]).
    A direct worst-case to expander-case self-reduction to a graph problem $\mathcal{A}$, is an algorithm that given any instance $G$ with $n$ vertices and $m$ edges, computes in $\tilde{O}(n+m)$ time a graph $G_{exp}:=(V_{exp},E_{exp})$ with the following guarantees:
    • $G_{exp}$ is an $\Omega(1)$-expander with high probability.
    • The blowup in the number of vertices and edges in $G_{exp}$ is bounded. Namely, $|V_{exp}|\leq K$ and $|E_{exp}|\leq M$ for some $K:=K(n,m)$ and $M:=M(n,m)$.
    • The solution $\mathcal{A}(G)$ can be computed from the solution $\mathcal{A}(G_{exp})$ in $\tilde{O}(m+n)$ time.
    Direct-WTERs can be used to show equivalence between the complexity of polynomial-time problems and their complexity on $\Omega(1)$-expanders. Namely, if problem $\mathcal{A}$ is a polynomial-time problem that admits a Direct-WTER, then $\Omega(1)$-expanders are worst-case instances of $\mathcal{A}$ (ignoring poly-logarithmic factors).
    The main contribution of Abboud and Wallheimer was to show that some fundamental problems, such as $k$-Clique Detection and Maximum Matching, admit simple Direct-WTERs. In particular, their Direct-WTERs make a graph an expander by employing a core gadget that augments it with $O(n)$ vertices and $\tilde{O}(n+m)$ random edges and then applies additional gadgets that control the solution. In particular, they obtain a near-linear blowup. Their results are surprising because such Direct-WTERs do not employ any of the heavy machinery that usually comes with expander decompositions, yet they output quantitatively better expanders: the outputs of Direct-WTERs are $\Omega(1)$-expanders (by definition), whereas expander decompositions can only produce $O(1/\log n)$-expanders (that are not as expanding) [39, 6].
    Furthermore, the simplicity of such Direct-WTERs leads to interesting and important messages to algorithm designers, as observed in [3]: The expander decomposition method is useless in the presence of Direct-WTERs because decomposing a graph into $o(1)$-expanders is meaningless when we can assume that the input graph is already an $\Omega(1)$-expander after a simple modification. This addressed (with a negative answer) a question that many researchers have wondered about as they looked for the next breakthrough to be obtained via the expander decomposition method:
    Question 2.
    Are expander decompositions the key to solving my problem?
This work.
Motivated by the appeal of the method of Direct-WTERs towards answering Question 1 and Question 2, our goal is to develop this theory further. Toward that, we address the two main limitations that were highlighted in [3]: (1) the Direct-WTERs are randomized whereas ED-WTERs are deterministic [16], and (2) the results are restricted to the Word-RAM model, whereas expander decompositions are popular tools in other models as well. In addition, we continue their line of work by providing Direct-WTERs to additional problems. Let us motivate these two topics before stating our results formally.
1.1 Deterministic Direct-WTERs
The randomized Direct-WTERs of Abboud and Wallheimer [3] prove that $\Omega(1)$-expanders are worst-case instances of many problems if we allow algorithms to be randomized. They leave us wondering if perhaps $\Omega(1)$-expanders are truly easier for deterministic algorithms. We remark that while it is believed that all algorithms can be derandomized by incurring a small polynomial blowup (as in $P=BPP$), it is far from clear that this blowup can be made $n^{o(1)}$ (see [15] for the state-of-the-art on such results). Can we provide a positive answer to Question 1 with respect to deterministic algorithms by designing deterministic Direct-WTERs?
Additional motivation comes from the hope of using the expander decomposition method in order to get breakthrough derandomization results, along the lines of Question 2, for problems where the current randomized algorithms are much faster than the current deterministic algorithms. Indeed, deterministic expander decompositions [16] have already played a major role in some of the most remarkable derandomization results of recent years, e.g. for Global Min-Cut [29, 37, 31, 30, 25].
Question 3.
Are expander decompositions the key to derandomizing my algorithm?
Similarly to the observation in [3], deterministic Direct-WTER also convey a message to algorithm designers: that the answer to the above question is negative, i.e., expander decompositions are useless for derandomizing the problem.
Motivated by this, our first result is a derandomization of the core gadget in [3], resulting in deterministic Direct-WTERs for various problems. In particular, we show that all problems admitting randomized Direct-WTERs in [3], and some additional problems, such as the Max-Cut problem, admit deterministic direct-WTERs.
Theorem 2.
The following problems admit deterministic Direct-WTERs: Maximum Matching, Minimum Vertex Cover, $k$-Clique Detection, $k$-Clique Counting, Max-Clique, Max-Cut, Minimum dominating Set, and $H$-Subgraph Detection ($m=\tilde{O}(m)$ and $H$ does not contain pendant vertices).
We provide formal definitions and an overview of all problems mentioned below in Appendix D.
An important feature of our deterministic core gadget is that it remains simple and efficient. Interestingly, this stands in contrast with other derandomization results in fine-grained complexity, which often tends to involve sophisticated methods and some slowdown (see, e.g.,  [11, 18]). As discussed earlier, simplicity is an important aspect of Direct-WTERs, not just because it makes them more accessible to the community, but also because it strengthens the message that expander decompositions become useless in the presence of Direct-WTERs. In Section 2, we provide an intuitive overview of how our new results are obtained by a modification to the core gadget of Abboud and Wallheimer [3]. Section 4 provides the construction itself and the full analysis.
Interestingly, our Direct-WTERs also improve upon the blowup in the number of added edges over the Direct-WTERs in [3], resulting in $\Omega(1)$-expanders with $O(n)$ vertices and $O(m+n)$ edges, whereas in [3], the expanders have $O(m+n\log n)$ edges. Note that this blowup is optimal since any expander is connected and, therefore, must contain $\Omega(n)$ edges. We also demonstrate that our core gadget preserves the following graph properties: (1) Bipartiteness-preserving; we can modify the core gadget so that if $G$ is bipartite, then so is the expander, and (2) Degree-preserving; if the maximum degree in $G$ is $\Delta$, then the maximum-degree in the expander is $2\Delta+O(1)$.
Remarks:
  1. 1.
    For the exponential-time problems in Theorem 2: Minimum Vertex Cover, Minimum Dominating Set, Max-Clique, and exact Max-Cut, the blowup in the number of vertices in the output graph must be subject to stronger restrictions than for polynomial-time problems. In particular, the blowup should be $n+o(n)$, to show that expanders are worst-case instances.333 Otherwise, an exponential speed-up on expanders does not necessarily translate to an exponential speed-up on general graphs. For such problems, we employ a generalized core gadget, which gives a tradeoff between the conductance and the blowup, resulting in Direct-WTERs providing, for every $0<\varepsilon\leq 1$, conductance $\Omega(\varepsilon)$ and blowup $\varepsilon n$.
  2. 2.
    For $k$-Clique Detection, we slightly improve the parameters over the Direct-WTERs given in [3] for this problem. In [3], the output is an $\Omega(1/k^{2})$-expander $G_{exp}$, with a blowup of $\Theta(nk)$ vertices and $\Theta(k^{2}m)$ edges, where each $k$-clique in $G$ corresponds to $k!$ $k$-cliques in $G_{exp}$. Our improved Direct-WTER produces an $\Omega(1)$-expander $G_{exp}$ with $O(n)$ vertices and $O(m+n)$ edges (even if $k=\omega(1)$), such that every $k$-clique in $G$ corresponds to $k+1$ $k$-cliques in $G_{exp}$. This enhancement makes our reduction more suitable for parameterized algorithms and larger (non-constant) values of $k$.
  3. 3.
    For the Max-Cut problem, there is an interesting related work on the approximation variant of the problem on expanders. A famous algorithm by Goemans and Williamson [20] obtains a $>0.878$-approximation for the maximum cut in general graphs, based on a Semidefinite Programming relaxation. In search of other, perhaps simpler methods for approximating the max-cut beyond the trivially obtained $1/2$-approximation [^4] Which follows from the fact that the maximum cut is at least $m/2$ in any graph with $m$ edges., Trevisan [42] posed the following question. Is there a combinatorial algorithm that achieves better than $1/2$-approximation?
    A positive answer to this question was given by Kale and Seshadhri [27], which remains the current-best combinatorial algorithm for this problem. A recent paper by Peng and Yoshida [36] addresses this question on expanders, providing a combinatorial algorithm for approximating the maximum cut on $\phi$-expanders. Namely, the authors provide an algorithm that given $\varepsilon$, computes a $(1/2+\varepsilon)$-approximation, subject to $\varepsilon=O(\phi^{2})$. Specifically, it computes a value $x$ such that $(1/2+\varepsilon)MC(G)\leq x\leq MC(G)$, where $MC(\cdot)$ denotes the cardinality of the maximum cut. Moreover, its running time is sublinear when $\phi$ is a constant. Our Direct-WTER for Max-Cut in Theorem 2, on the other hand, is a reduction that given $G$ and $0<\phi\leq 1$, outputs an $\Omega(\phi)$-expander $G_{exp}$, and the maximum cut in $G_{exp}$ is $MC(G_{exp})\leq(1+4\phi)MC(G)$, assuming the input graph is not too sparse [^5] More precisely, this holds if in $G$ there are $m>Cn$ edges for some constant $C$..
    Can we apply the Direct-WTER and then use the algorithm of Peng and Yoshida to get a combinatorial algorithm that $(1/2+\varepsilon)$-approximates the maximum cut in general graphs? Perhaps we may even improve upon the algorithm by Kale and Seshadhri, as both the Direct-WTER and the algorithm of Peng and Yoshida are very efficient. However, the approximate value we get from this approach is $(1/2+\varepsilon)MC(G)\leq x\leq(1+4\phi)MC(G)$, or equivalently $(1/2+\varepsilon)/(1+4\phi)MC(G)\leq x\leq MC(G)$. For this approximation ratio to be larger than $1/2$, we need to pick $\varepsilon>2\phi$, but recall the constraint $\varepsilon\leq O(\phi^{2})$. Hence, this approach fails. This is not so surprising, since our Direct-WTER is quite elementary so we do not expect to make use of it as a subroutine inside another algorithm. Instead, we can interpret this result as a limitation to algorithms for Max-Cut on $\phi$-expanders; that one cannot obtain a $(1/2+\varepsilon)$-approximation for some $22\phi<\varepsilon\leq 1/$ 2 unless this highly unlikely approach works.
1.2 Direct-WTERs in the Fully Dynamic setting
Before this work, Direct-WTERs were limited to the (randomized) Word-RAM model, whereas ED-WTERs could address many other models. One particular area in which expanders, expander decomposition, and derandomization are important subjects is the area of dynamic graph algorithms. Let us focus on the Fully Dynamic model of computation, where the goal is to maintain the solution of a problem in a graph undergoing edge updates, i.e., edge insertions and deletions. Dedicated tools have been developed for maintaining an expander decomposition in this model and subsequently achieved major breakthroughs (e.g., [33]). Notably, derandomization is a central concern in the dynamic setting because deterministic algorithms are essentially the only ones that work against adaptive adversaries (see, e.g., [9]). For these reasons, the three main questions outlined above are particularly interesting in this model. Our first question, in this context, is whether Direct-WTERs can be adapted to the Fully Dynamic setting.
An important related work is a recent paper by Henzinger, Paz, and Sricharan [26] (abbreviated as HPS), who initiated the study of Question 1 in the dynamic model, regarding the complexity of fundamental problems on dynamic expanders. A dynamic expander is a dynamic graph that undergoes edge updates but remains an $\Omega(1)$-expander at any point in time. They adapt lower bound proofs from fine-grained complexity, so that they hold even on dynamic expanders. Their techniques differ from the self-reduction approach of Direct-WTERs and correspond to the first of the three methods outlined above to answer Question 1. In particular, they base their results on the Online Matrix Vector (OMv) conjecture, which was introduced by Henzinger et al. [24] to prove lower bounds for various dynamic problems. HPS obtained their results by adapting these lower bound proofs to the case of constant-degree expanders, proving that dynamic expanders whose maximum degree remains bounded by a constant are OMv-hard. The problems they consider in this context are Maximum Matching, Densest Subgraph, and $st$-Shortest Path (abbreviated as $st$-SP). This leaves us wondering with the following questions:
  1. 1.
    Are expanders in higher density regimes, whose maximum degree is not bounded by a constant, also OMv-hard instances?
  2. 2.
    Can the techniques that are used in the static setting to prove Theorem 2 contribute to this study by, e.g., providing a simpler, or alternative method to prove that certain problems remain OMv-hard on expanders?
Our second main result is an adaptation of the deterministic core gadget to the dynamic setting, resulting in a deterministic, dynamic algorithm for maintaining a dynamic $\Omega(1)$-expander, whose running time is amortized $\tilde{O}(1)$ per edge update. Subsequently, we show Dynamic Direct-WTERs (abbreviated as DD-WTERs) to various problems, [^6] Namely, to all problems in the previous theorem, except that we do not discuss the exponential-time problems in the dynamic model in this work. thus proving that $\Omega(1)$-expanders are worst-case instances and that the expander decomposition method is useless against them. We present and discuss the formal definition of DD-WTERs in Appendix B.
Theorem 3.
The following problems admit a DD-WTER: Maximum Matching, Bipartite Perfect Matching, Densest Subgraph (in graphs with $m>42n$ edges), $k$-Clique Detection, $k$-Clique Counting, and $H$-Subgraph Detection (where $m=\tilde{O}(n)$ and $H$ does not contain pendant vertices).
In addition, our results also have some interesting implications related to the work of HPS and the questions above.
  1. 1.
    For the Densest Subgraph problem, Henzinger et al. [24] prove a lower bound of $n^{1/3-o(1)}$ per update under OMv for general graphs. By modifying their reduction, HPS were able to show a weaker $n^{1/4-o(1)}$ lower bound for expanders.777 We remark that the proof in Henzinger et al. [24, Corollary 3.26] gives an $n^{1/3-o(1)}$ lower bound under OMv, while the introductions of both [24] and HPS [26] mention, erroneously, an $n^{1/2-o(1)}$ lower bound.
    As a consequence of our direct DD-WTER, we conclude that the $n^{1/3-o(1)}$ lower bound for general graphs also holds for expanders. One subtlety towards this result is that the reduction of Henzinger et al. produces very sparse graphs with $m\leq 2n$ edges while our DD-WTER assumes that $m>42n$. In Appendix A, we discuss how to modify the original reduction so that denser graphs are produced. Another strength of the DD-WTER compared to HPS is that it is a self-reduction, hence it does not depend on the OMv Conjecture to get a lower bound (at least when the $m>42n$ assumption is made) [^8] We remark that the limitation of $m>42n$ in this approach is due to the fact that to augment a graph to become an expander requires at least adding some amount of edges, which unavoidably affects the densest subgraph..
  2. 2.
    For the Maximum Matching problem on graphs of maximum degree $O(n^{t})$, for any $0\leq t\leq 1$, we show that this problem is OMv-hard on $\Omega(1)$-expanders as well. This improves upon HPS, who only prove it for the constant-degree case (i.e. $t=0$). Our result is obtained by combining a lower bound proof from HPS for graphs of maximum degree $O(n^{t})$, with our degree-preserving DD-WTER for Maximum Matching.
    Corollary 4.
    For any $0\leq\varepsilon,t\leq 1$ and any constant $\varepsilon>0$, there is no dynamic algorithm for maintaining a maximum matching on $\Omega(1)$-expanders with maximum degree $O(n^{t})$, with amortized $O(n^{(1+t)/2-\varepsilon})$ update time and $O(n^{1+t-\varepsilon})$ query time, unless the OMv Conjecture is false.
  3. 3.
    We demonstrate that our core gadget is useful even outside the context of self-reductions by showing a DD-WTER for Graphical OMv, an equivalent graph formulation of the OMv problem. This implies that we can lift OMv-based lower bounds to $\Omega(1)$-expanders for many problems, in particular, for problems such as Maximum Matching and $st$-SP, which HPS considered. In particular, we demonstrate the power of this technique by proving that $st$-SP is OMv-hard on $\Omega(1)$-expanders. [^9] Note that in comparison with HPS, we do not prove hardness for graphs of constant degree, which is outside the scope of our paper.
    Proposition 5 ($st$-SP is OMv-hard on $\Omega(1)$-expanders).
    For any $\varepsilon>0$, there is no dynamic algorithm for the dynamic $st$-SP problem on $\Omega(1)$-expanders, with polynomial preprocessing time, $O(m^{1/2-\varepsilon})$ update time, and $O(m^{1-\varepsilon})$ query time, assuming the OMv Conjecture.
Finally, let us discuss another aspect of how our work and HPS differ. One of the main motivations for our work is to address Questions 2 and 3 about the applicability of expander decompositions in algorithms, whereas HPS main motivation is to gain a better understanding of the complexity of dynamic problems on various graph families, including expanders, along the lines of Question 1. As we have seen, DD-WTERs imply that expander decompositions are useless because they show equivalence between the worst-case and the expander-case complexities. We remark that the results of HPS imply the same, assuming their obtained lower bounds are tight. Hence, it implies a conditional answer to these questions, whereas self-reductions imply an unconditional one.
1.3 Expanders in distributed models.
The above results demonstrate that the impact of Direct-WTERs goes beyond the classic Word-RAM model of computation. A natural continuation of this is to apply these techniques to more models of computation, where expanders, expander decompositions, and derandomization are important subjects. In Section 8, we discuss the applicability of Direct-WTERs in distributed models of computation. We show that while there are limitations in adapting Direct-WTERs to CONGEST, it is possible to do so in CONGESTED-CLIQUE and MPC.
Roadmap.
A technical overview is given in Section 2, where we also explain the differences compared to [3]. Then, after some preliminaries in Section 3, we provide the full details of our derandomized core gadget in Section 4. Section 5 presents variants of the core gadget, including a fully dynamic adaptation. Section 6 presents deterministic Direct-WTERs for Max-Cut, Densest Subgraph, and Graphical OMv. In addition, Section 7 presents a dernadomization and dynamization of all the Direct-WTERs that appeared in [3], and additional Direct-WTERs for related problems. Section 8 addresses Direct-WTERs in distributed models of computation. Appendix B contains a formal definition of DD-WTERs.
2 Technical Overview
In this section, we summarize the core gadget of Abboud and Wallheimer [3] (henceforth, AW) that is used in all their Direct-WTERs, and then present our modification. The techniques in AW’s work are based on a simple core gadget that takes a graph $G$ as input and augments it with vertices and edges to output an expander $G_{exp}$ containing $G$ as an induced subgraph. Roughly, their construction boils down to the following procedure.
AW’s core gadget.
Given a graph $G=(V,E)$, add a set $U$ of $n$ vertices called expansion layer. Then, for every vertex $v\in V$, take a sample of $O(\log n)+\deg_{G}(v)$ vertices in $U$ and make them neighbors of $v$. Clearly, the size of this graph is $O(n\log n+m)$. In addition, they prove that it is an $\Omega(1)$-expander with high probability by showing that the probability that a cut in $G_{exp}$ ends up being sparse is very low. Hence, the result follows the union bound. [^10] We remark that their construction also contains some additional $O(n\log n)$ edges, which are added deterministically, but careful inspection shows that they are not needed. The time complexity to construct the expander is $\tilde{O}(n+m)$ time. However, there is no guarantee that the problem’s solution to $G$ can be computed easily from the solutions to $G_{exp}$. To this end, AW provide additional, mostly simple gadgets to control the solution in $G_{exp}$ in a predictable manner while preserving the conductance of $G_{exp}$ up to a constant factor.
Derandomizing the core gadget.
The first limitation of AW’s approach that we aim to resolve in this paper is their use of randomness. To do so, one could attempt to introduce an explicit, $d$-regular bipartite $\Omega(1)$-expander $X$ between $V$ and $U$. However, which degree $d$ should we choose? For starters, we consider picking some constant $d\geq 3$. However, this approach might fail if $G$ contains cuts with many internal edges and few outgoing edges. For instance, if $G$ includes a cut $S$ of $\sqrt{n}$ vertices that form an isolated $\sqrt{n}$-clique, then in $G_{exp}$ we have $vol(S)=\Theta(n)$, while the number of outgoing edges added by the expander is $O(|S|d)=O(\sqrt{n})$, thus $\phi(S)=O(1/\sqrt{n})$. On the opposite end, we could select a complete bipartite graph with $\Theta(n^{2})$ edges. While this will address the problem and indeed result in an $\Omega(1)$-expander, this approach is undesirable due to the excessive number of added edges. Hence, the most natural strategy to consider is incorporating a $d$-regular expander for $d=\lceil\frac{m}{n}\rceil$. To be more precise, since explicit constructions exist only for $d\geq 3$, the approach is to pick $d=\lceil\frac{m}{n}\rceil+3$ (as it could be the case that, for example, $m=0$). While the number of added edges is now $O(m+n)$, this approach is slightly naïve, as it does not even address the above problem of isolated $\sqrt{n}$-clique in a graph with $O(n)$ edges. The crux of the issue lies in the fact that $G$ is not regular. If $G$ were $d$-regular, then the inclusion of a $d$-regular expander would resolve our problem. One can look for expanders with a certain degree sequence, matching the one of $G$, but constructing those seems challenging.
The starting point for our modification is the observation that the randomness in AW’s constructions is coming to accomplish two things at once: first and foremost, choosing random edges breaks structure and creates a “random like” graph. However, second, and less obviously, it is a way to achieve a certain “balanced allocation” of neighbors in $U$, resulting in nearly uniform degrees in $U$, which is crucial for their analysis to work. This leads to our modification, in which we substitute their single expansion layer $U$ with two layers: one for balanced allocation denoted $L$, connected to $V$ using a load-balancing algorithm, ensuring that the degrees in this layer are about $\frac{2m}{n}$, and another layer for expansion denoted $R$, connected to $L$ using the edges of a bipartite, $(\frac{2m}{n}+3)$-regular $\Omega(1)$-expander. A random graph can accomplish the construction of the first layer, as AW did, but it could also be accomplished in any other way, such as the standard Round-Robin algorithm without any randomness. Then, all we have to do is use an explicit construction (described in Appendix C) of a regular $\Omega(1)$-expander. This is the whole core gadget in Section 4. Using it, we follow the steps of AW and add various gadgets to $G_{exp}$, such that the solution for $G$ can be retrieved from the solution for $G_{exp}$.
Dynamizing the core gadget.
The second limitation of AW’s approach, which we aim to resolve, is that it focuses solely on the static model of computation. To dynamize our deterministic core gadget, we employ a lazy-greedy approach as follows. As edges are inserted and deleted from $G$, we lazily wait until some local or global imbalance event occurs. The local imbalance event is when the degree of a vertex in $G$ doubles itself. In this case, we add more edges from this vertex to $L$ using a (deterministic) greedy balanced allocation strategy. We prove that this greedy strategy ensures that the degrees in $L$ remain within a constant factor from each other for a long time, which is a sufficient condition for the graph to remain an $\Omega(1)$-expander. The global imbalance events are as follows. One event is when the degrees in $L$ become overly imbalanced, in which case we apply a costly rebalance procedure using Round-Robin. Another global imbalance event is when the number of edges in $G$ differs too much from the number of edges in $G_{exp}$, in which case we apply a costly reconstruction procedure using the static core gadget. This approach results in an amortized slowdown of $O(1)$ time per operation.
3 Preliminaries
Let $G=(V,E)$ be an undirected simple graph. Throughout the paper, we use $n$ and $m$ to denote the number of edges in vertices in $G$, respectively. Denote th neighborhood of $v$ by $N(v):=\{u\in V\mid(v,u)\in E\}$ and the degree of $v$ by $\deg(v)$. We say that $G$ is $d$-regular if $\deg(v)=d$ for all $v\in V$. A vertex $v$ is called a pendant vertex if $\deg(v)=1$. The set of all edges with one endpoint in $S\subseteq V$ and another endpoint in $T\subseteq V$ is denoted by $E(S,T):=\{(u,v)\in E\mid u\in S,v\in T\}$, and its cardinality is denoted by $e(S,T):=|E(S,T)|$. We call $E(S,V\setminus S)$ the out-going edges of $S$. We employ subscripts to indicate which graphs we refer to when it is not clear from the context. For instance, $\deg_{H}(v)$ denotes the degree of vertex $v$ in a graph $H$.
Conductance.
Let $S\subseteq V$ be a cut. The volume of $S$ is defined as $vol(S):=\sum_{v\in S}\deg(v)$. The conductance of $S$ is defined as $\phi(S):=e(S,V\setminus S)/\min(vol(S),vol(V\setminus S))$. If $\min(vol(S),vol(V\setminus S))=0$, we define $\phi(S)=0$. The conductance of the entire graph $G$ is defined as $\phi_{G}:=\min_{\emptyset\neq S\subseteq V}\phi(S)$. Throughout this paper, unless explicitly indicated otherwise, we adhere to the following definition of expander graphs, in which $0\leq\phi\leq 1$.
Definition 6.
$G$ is a $\phi$-expander if $\phi_{G}\geq\phi$.
This definition aligns with the one commonly used in the literature about expander decompositions.
Edge-expansion.
Another notion of expansion that we use in our proofs is edge expansion.
Definition 7.
The edge expansion of $G$ is:
$h_{G}:=\min_{\emptyset\neq S\subseteq V,|S|\leq n/2}\frac{e(S,V\setminus S)}{|S|}.$
We will say that $G$ is an $h$-edge expander if $h_{G}\geq h$.
A known fact states that conductance and edge expansion are interchangeable in regular graphs.
Fact 8.
If $G$ is a $d$-regular $h$-edge expander, then it is also a $\frac{h}{d}$-expander. In particular, if $G$ is an $\Omega(d)$-edge expander, then it is also an $\Omega(1)$-expander.
4 The Core Gadget
In this section, we present a derandomized core gadget on which we base our results. The core gadget is a deterministic algorithm that takes a graph $G$ and augments it with new vertices and edges to output an $\Omega(1)$-expander $G_{exp}$ in $\tilde{O}(m+n)$ time. We will utilize an explicit construction of $d$-regular, bipartite, $\phi d$-edge expanders on $2N$ vertices for some constant $\phi>0$. Assume that given any $d\geq 3$, and sufficiently large $N$, we can construct such graph in $\tilde{O}(Nd)$ time. See Appendix C for details about the construction.
4.1 The core gadget
Given a graph $G$, augment $G$ with an initially empty bipartite graph featuring $N$ vertices on each side, denoted $L:=\{x_{1},x_{2},\ldots,x_{N}\}$ and $R:=\{y_{1},y_{2},\ldots,y_{N}\}$. $N$ is a parameter that can vary depending on the application, but for the most basic construction, we set it to $N=(1+o(1))n$. The gadget consists of the following two-step construction:
  1. 1.
    (Degree balancing) For every $v\in V$ add, using a Round-Robin algorithm, $\deg_{G}(v)+3$ neighbors in $L$. Namely, follow a circular order over $L$ to pick neighbors one at a time. Assuming $N\geq n+2$, every vertex is connected to $\deg_{G}(v)+3$ distinct neighbors in $L$ without wrapping around. The Round-Robin algorithm guarantees that the degrees within $L$ are almost balanced. Namely, since the total number of edges that we add in this step is $\sum_{v\in V}(\deg_{G}(v)+3)=2m+3n$, the degree of every vertex in $L$ is either $\lfloor\frac{2m+3n}{N}\rfloor$ or $\lceil\frac{2m+3n}{N}\rceil$.
  2. 2.
    (Expander construction) Set $d:=\lceil\frac{2m+3n}{N}\rceil$. Let $X$ be a bipartite, $d$-regular, $\phi d$-edge expander on $L\cup R$, constructed using the algorithm from Appendix C.
Denote the resulting graph by $G_{exp}:=(V_{exp},E_{exp})$, illustrated in Figure 1. Note that the running time of this gadget is $\tilde{O}(m+n+dN)=\tilde{O}(m+n)$.
Figure 1: The core gadget augments $G$ with $O(m+n)$ edges connected to a bipartite, $d$-regular, $\phi d$-edge expander, resulting in an $\Omega(1)$-expander $G_{exp}$.
For clarity, we henceforth omit the subscript $G_{exp}$ from $e_{G_{exp}}(\cdot,\cdot)$, and keep the subscript $G$ in $e_{G}(\cdot,\cdot)$. Similarly, we do for $E(\cdot,\cdot)$, $\deg(\cdot)$, and $vol(\cdot)$. Let us now prove the main lemma.
Lemma 9.
The graph $G_{exp}$ is an $\Omega(1)$-expander.
On the high level, we split into two cases. In the first case, we deal with cuts whose volume is mostly concentrated in $X$ and show that there must be many outgoing edges because $X$ is an expander, and the degrees in $L$ are bounded. In the second case, we deal with cuts whose volume is mostly concentrated in $G$ and show that there must be many outgoing $V$-to-$L$ edges in this case. See Figure 2. In more detail:
Proof.
To prove that $G_{exp}$ is an $\Omega(1)$-expander, consider a non-empty cut $S\subseteq V_{exp}$. We need to show that $\min(vol(S),vol(V_{exp}\setminus S))>0$, and that $e(S,V_{exp}\setminus S)\geq\phi\cdot\min(vol(S),vol(V_{exp}\setminus S))$ for some constant $\phi>0$. It is clear that $\min(vol(S),vol(V_{exp}\setminus S))>0$ because all vertices have a degree of at least $3$. Let $S_{V}:=S\cap V$, $S_{L}:=S\cap L$, and $S_{R}:=S\cap R$. Assume w.l.o.g. that $|S_{L}\cup S_{R}|\leq N$, as it holds either for $S$ or its complement. Note that we have $vol(S)=vol(S_{V})+vol(S_{L}\cup S_{R})$, and consider the following cases:
  • $vol(S_{V})\leq 4vol(S_{L}\cup S_{R})$. In this case, $vol(S)\leq 5vol(S_{L}\cup S_{R})$, and since every $x\in L$ has degree $2d$ or $12d-$ 1, and every $y\in R$ has degree $d$, then $5vol(S_{L}\cup S_{R})\leq 10d|S_{L}\cup S_{R}|$. Since $X$ is a $d$-regular, $\phi d$-edge expander on $2N$ vertices, and $|S_{L}\cup S_{R}|\leq N$, there is some constant $\phi>0$ such that $e(S_{L}\cup S_{R},(L\cup R)\setminus(S_{L}\cup S_{R}))\geq\phi d|S_{L}\cup S_{R}|$. Hence, $e(S,V_{exp}\setminus S)\geq\phi d|S_{L}\cup S_{R}|\geq\phi vol(S)/10$, and we are done.
  • $vol(S_{V})>4vol(S_{L}\cup S_{R})$. In this case, $vol(S)<5vol(S_{V})/4$. We show that there are many $S_{V}$-to-$(L\setminus S_{L})$ edges. Note that $e(S_{V},L)\geq vol(S_{V})/2$, because for every vertex $v\in V$, there are more $v$-to-$L$ edges than $v$-to-$V$ edges. Hence:
    $\begin{split}&e(S_{V},L\setminus S_{L})=e(S_{V},L)-e(S_{V},S_{L})\geq\frac{vol(S_{V})}{2}-vol(S_{L}\cup S_{R})>\\ &\frac{vol(S_{V})}{2}-\frac{vol(S_{V})}{4}=\frac{vol(S_{V})}{4}\geq vol(S)/5,\end{split}$ (1)
    and we are done.
Figure 2: An illustration for the proof of Lemma 9. In the first case, we deal with cuts whose volume is mostly concentrated in $X$ and show that there are many short-dashed edges because $X$ is an expander, and the degrees in $L$ are bounded. In the second case, we deal with cuts whose volume is mostly concentrated in $G$ and show that, in this case, there must be many long-dashed edges.
We remark that the blowup of the core gadget, i.e., the number of added vertices and edges, is $2N=O(n)$ vertices and $O(Nd)=O(m+n)$ edges. Furthermore, if $G$ is a graph of maximum degree $\Delta$, then $G_{exp}$ is a graph of maximum degree $32\Delta+$ 3.
5 Variants of the Core Gadget
In this section, we discuss a few variants of the core gadget and how to adapt it to the fully dynamic setting.
5.1 Robust core gadget
The first variant we consider is a generalization of the core gadget, which allows greater variations in the degrees of the vertices in $G_{exp}$. This generalization will be useful in the analysis of various Direct-WTERs both in the static and in the dynamic setting. Namely, the next lemma shows that changing the degrees by a constant factor reduces the conductance by, at most, a constant factor.
Lemma 10.
For every $0<\varepsilon\leq 1$, $\alpha\geq 1$, and an integer $d_{X}\geq 3$, consider the following generalization of the core gadget:
  1. (1)
    Every $v\in V$ has at least $\varepsilon\deg_{G}(v)+1$ neighbors in $L$.
  2. (2)
    $X$ is a $d_{X}$-regular, $\phi d_{X}$-edge expander, for some some constant $\phi>0$.
  3. (3)
    The degrees of all the vertices in $L$ are within $[d_{X},\alpha d_{X}]$.
Then $G_{exp}$ is an $\phi\varepsilon/(5\alpha)$-expander.
Proof.
The proof is obtained by adapting the proof of Lemma 9 to this generalization. For brevity, we only discuss its differences from the proof of Lemma 9. Note that $vol(S)>0$ for every $S\subseteq V_{exp}$ because all vertices have positive degrees. Let us adapt the two cases as follows:
  • $vol(S_{V})\leq\frac{4}{\varepsilon}vol(S_{L}\cup S_{R})$. In this case, $vol(S)\leq(1+\frac{4}{\varepsilon})vol(S_{L}\cup S_{R})\leq(1+\frac{4}{\varepsilon})\alpha d_{X}|S_{L}\cup S_{R}|$. Therefore, since $X$ is a $\phi d_{X}$-edge expander, we get:
    $\begin{split}&e(S_{L}\cup S_{R},(L\cup R)\setminus(S_{L}\cup S_{R}))\geq\phi d_{X}|S_{L}\cup S_{R}|\geq\frac{\phi}{\alpha(1+\frac{4}{\varepsilon})}vol(S)\geq\frac{\phi\varepsilon}{5\alpha}vol(S),\end{split}$
    and we are done.
  • $vol(S_{V})>\frac{4}{\varepsilon}vol(S_{L}\cup S_{R})$. In this case, $vol(S)\leq(1+\frac{\varepsilon}{4})vol(S_{V})$. Note that $e(S_{V},L)\geq\frac{\varepsilon}{\varepsilon+1}vol(S_{V})\geq\frac{\varepsilon}{2}vol(S_{V})$, because every $v\in V$ has $\deg_{G}(v)$ neighbors in $V$ and at least $\varepsilon\deg_{G}(v)$ neighbors in $L$. Hence:
    $\begin{split}&e(S_{V},L\setminus S_{L})=e(S_{V},L)-e(S_{V},S_{L})\geq\frac{\varepsilon}{2}vol(S_{V})-vol(S_{L}\cup S_{R})>\\ &\frac{\varepsilon}{2}vol(S_{V})-\frac{\varepsilon}{4}vol(S_{V})=\frac{\varepsilon}{4}vol(S_{V})\geq\varepsilon vol(S)/5,\end{split}$ (2)
    and we are done.
5.2 Fully-dynamic core gadget
The above core gadget is not suitable for the fully dynamic setting, where $G$ undergoes edge insertions and deletions. Let us discuss the difficulties in dynamizing it and our high-level approach. Two issues arise in this setting. The first one is that when the degree of a vertex $v\in V$ increases due to edge insertions, we need to allocate to $v$ additional neighbors from $L$ while preserving balance in the degrees in $L$. One attempt to solve this issue is to pick a minimum-degree vertex in $L$, which can be computed quickly using a priority queue. While this ensures balanced degrees in $L$, there is a subtle issue that the minimum-degree vertex in $L$ might already be a neighbor of $v$; hence we can not add another edge to it. [^11] In the multigraph setting, this would not become an issue, as we can connect two vertices with multiple edges. However, our focus in this paper is on the more restricted family of simple graphs and simple expanders.
Therefore, we slightly modify this heuristic, and instead, our suggested greedy approach is to compute the successor of the minimum-degree vertex repeatedly, until we find a minimum-degree vertex in $L\setminus N(v)$, and then make it a neighbor of $v$. The required computation here is proportional to the degree of $v$, which might not be constant. Hence, we employ lazy updates; we wait until the degree of $v$ doubles itself before computing a new batch of neighbors in $L$ so that the amortized cost will be constant. One might notice, however, that the minimum-degree vertex in $L\setminus N(v)$ is not necessarily a minimum-degree vertex in $L$, so it needs to be clarified that the degrees in $L$ remain balanced. Nonetheless, we manage to show that the degrees in $L$ become overly imbalanced only after $\Omega(m+n)$ edge insertions. In this case, there is enough credit to reconstruct $E(V,L)$ using the Round-Robin algorithm, resulting in a total amortized time $O(1)$ per edge insertion. In general, edge deletions do not require further computation. However, too many edge deletions will make $X$ too dense compared to $G$, resulting in a large blowup. This is the second core issue.
The second issue is to maintain an appropriate degree of $X$. Recall that the degree of $X$ is picked to a value that is essentially the average degree of $G$. Therefore, if $G$ undergoes too many edge insertions, it may result in $X$ being too sparse, and the conductance of $G^{exp}$ becomes small. On the other hand, if $G$ undergoes too many edge deletions, then $X$ becomes too dense, and the blowup becomes large. To resolve this issue, we utilize the lazy updates approach: we recompute $X$ if its degree is too large or too small. This adds additional amortized cost of $O(1)$. Let us now provide the full details of the dynamized core gadget.
Our algorithm makes use of three procedures: (1) $\tt{UPDATE}$, which computes and adds a batch of edges from a vertex in $G$ to $L$, (2) $\tt{BALANCE}$, which rebalances the degrees in $L$ using Round-Robin, and (3) $\tt{RECOMPUTE}$, which recomputes the graph using the static core gadget. In this section, we use $G^{exp}$ to denote a dynamic $\Omega(1)$-expander output by the dynamized core gadget. We use subscripts to indicate the state of a dynamic graph at a certain time, e.g., $G_{t}$ is the dynamic graph $G$ at time $t$. We begin by describing the algorithm, then the three procedures, and lastly, we show the analysis of the algorithm.
Preprocessing
In the preprocessing step, given $G_{0}$, apply the static core gadget to construct an $\Omega(1)$-expander $G_{0}^{exp}$. We store the vertices of $L$ sorted according to their degrees, in a data structure supporting updates and successor queries in $O(1)$ time. [^12] Naïvely this would take $\tilde{O}(1)$ time using standard data structures, but it can be optimized to $O(1)$ since the degrees are integers in the range $[1,2N]$, and our updates only increase or decrease the degree of a vertex by $1$. We omit the details for the sake of brevity.
Edge insertions and deletions
For every insertion of an edge $uv$ to $G$, begin by inserting $uv$ to $G^{exp}$. Denote by $m_{t}$ the number of edges in $G$ at the last time we applied the $\tt{RECOMPUTE}$ procedure (or $m_{0}$ if we did not apply it). If $m\geq 2m_{t}+n$, apply the $\tt{RECOMPUTE}$ procedure to recompute the graph using the static core gadget and finish.
Otherwise, let us describe the process we apply to $v$ and similarly do to $u$. Denote by $\deg_{L}(v)$ the number of neighbors that $v$ has in $L$, i.e., $\deg_{L}(v)=|N(v)\cap L|$.
  1. (a)
    If $\deg_{G}(v)<2\deg_{L}(v)$, finish. Otherwise, apply the $\tt{UPDATE}$ procedure to $v$ to add a batch of neighbors to $v$, after which we have $\deg_{L}(v)=\deg_{G}(v)+3$.
  2. (b)
    Check if the degrees in $L$ became unbalanced, namely, if $\Delta_{L}\geq 2\delta_{L}$, where $\Delta_{L}$ and $\delta_{L}$ are the maximum and minimum-degree vertices in $L$, respectively. If so, apply the $\tt{BALANCE}$ procedure, after which $\Delta_{L}\in\{\delta_{L},\delta_{L}+1\}$.
For every deletion of an edge $uv$ from $G$, delete $uv$ from $G^{exp}$. Then, if $n\leq m\leq 0.5m_{t}$, apply the $\tt{RECOMPUTE}$ procedure to recompute $X$.
Let us now describe the three procedures used above.
  • $\tt{UPDATE}$ Let $k:=\deg_{L}(v)$. Compute $k+3$ minimum-degree vertices in $L\setminus N(v)$, denoted $x_{1},x_{2},\ldots,x_{k+3}$, by repeatedly making successor queries to $L$ and skipping vertices which belong to $N(v)$. For every $x_{i}$, insert an edge $vx_{i}$ to $G^{exp}$. Note that now we have $\deg_{L}(v)=\deg_{G}(v)+3$.
  • $\tt{BALANCE}$ Compute, using Round-Robin, a new set of $V$-to-$L$ edges, denoted $R$, and then replace $E(V,L)$ with $R$. Specifically, insert the edges of $R$ in the order given by the Round-Robin algorithm before removing $E(V,L)\setminus R$, to ensure that the degrees do not vary too much in the intermediate graphs.
  • $\tt{RECOMPUTE}$. Apply the static core gadget to compute a set of edges $R$, and $X^{\prime}$, where $R$ is the set of $V$-to-$L$ edges as before, and $X^{\prime}$ is the expander on $L\cup R$. In particular, $X^{\prime}$ is a $d_{X^{\prime}}$-regular, $\Omega(d_{X^{\prime}})$-edge expander for $d_{X^{\prime}}=\lceil\frac{2m+3n}{N}\rceil$. Replace $X$ with $X^{\prime}$ by first inserting the edges of $X^{\prime}$, and then removing the leftover edges of $X$ which do not belong to $X^{\prime}$. Then insert the edges of $R$ and remove $E(V,L)$, as we did above. We think of $X^{\prime}$ as being the “active” expander at every intermediate step of this procedure.
Now, let us analyze the algorithm, starting with the correctness.
Proposition 11.
Every intermediate graph output by the dynamic core gadget has blowup $O(m+n)$ and conductance $\Omega(1)$.
Proof.
The blowup is clearly $O(m+n)$ because at every intermediate step, we add at most $|R|+|e(X^{\prime})|=O(m+n)$ edges. Let us prove that the conductance is $\Omega(1)$. Observe that:
  • At any stage, we have $\deg_{L}(v)\geq\max(3,\deg_{G}(v)/2)$, which implies that $\deg_{L}(v)\geq\deg_{G}(v)/5+1$.
  • At any stage of $\tt{RECOMPUTE}$, the edge-expansion of the graph is at least $\phi d_{X^{\prime}}/4$ because the degree $X$ (the expander we discard) is at least $\phi d_{X^{\prime}}/4$. Note that adding edges to an edge expander does not ruin edge expansion.
  • At any stage, the degrees in $L$ are within $[d_{X},4d_{X}]$ for some constant $\alpha$. This follows since $\delta_{L}\geq d_{X}$ and $\Delta_{L}\leq 4\delta_{L}$.
Therefore, by Lemma 10, the graph is an $\Omega(1)$-expander. ∎
Now, let us analyze the performance of the core gadget.
Proposition 12.
The preprocessing time of the core gadget is $\tilde{O}(m+n)$, and the amortized update time is $O(1)$. In particular, for every sequence of $M$ updates to $G$, it computes in $\tilde{O}(m+n)+O(M)$ time a sequence of $O(M)$ updates to $G_{exp}$.
Proof.
We prove this proposition for each of the above procedures separately that their amortized cost is $O(1)$:
  • $\tt{UPDATE}$. The cost of this step is bounded by $O(k)$ (where $k=\deg_{L}(v)$) because there are at most $k+(k+3)$ successor queries to the data structure, and at most $k+3$ edge insertions to $G^{exp}$. Note that $k$ might not be a constant. However, since $\deg_{G}(v)=2k$ and $\deg_{L}(v)=k$, and after every call to $\tt{UPDATE(v)}$, we have $\deg_{L}(v)=\deg_{G}(v)+3$, there are $k+3$ edge insertions of edges incident to $v$ that we can charge to: the edges incident to $v$ that were inserted to $G$ since the last call to $\tt{UPDATE(v)}$. We charge the cost of UPDATE to them. Since there are at least $k$ many edges, and the cost of UPDATE is $O(k)$, the amortized cost is $O(1)$.
  • BALANCE. The cost of BALANCE is $O(m+n)$. To bound the amortized cost, we will show that every call to BALANCE is preceded by $\Omega(n+m)$ edge insertions to $G$. To see why, consider a graph obtained from a call to RECOMPUTE (or preprocessing), in which for every vertex $v\in V$, we have $\deg_{L}(v)=\deg_{G}(v)+3$. We claim that during edge insertions that follow, every insertion causing an increase in $\Delta_{L}$ (through UPDATE), except the first one, is preceded by at least $N/4$ edge insertions to $G$. To see why, let us denote by $N_{i}$ the number of vertices of degree $i$ in $L$. Consider a call to $\tt{UPDATE(v)}$ that increases $\Delta_{L}$ or $N_{\Delta_{L}}$ (or both). As we observed before, there are at least $k+3=\deg_{L}(v)+3$ edges we can charge to. We will either charge to those edges or reserve them to be charged later. Namely:
    • For every increase in $N_{\Delta_{L}}$: we reserve one uncharged edge to be charged later. Since UPDATE can increase $N_{\Delta_{L}}$ by at most $k+3$, the number of uncharged edges is sufficiently large. Hence, at any point in time, there is a reserve of $N_{\Delta_{L}}$ uncharged edges in the system.
    • If UPDATE increases $\Delta_{L}$: it means that $N-k-N_{\Delta_{L}}<k+3$, because the number of neighbors added to $v$, $k+3$, is greater than the number of vertices in $L\setminus N(v)$ that have degree $<\Delta_{L}$, and there are at least $N-k-N_{\Delta_{L}}$ such vertices. Now, if $N_{\Delta_{L}}<N/2$, then $N/4\leq k$, so we charge the increase in $\Delta_{L}$ to $k\geq N/4$ edge insertions. On the other hand, if $N_{\Delta_{L}}\geq N/2$, then we have a reserve of $\geq N/2$ uncharged edges. We charge the increase in $\Delta_{L}$ to this reserve.
    Therefore, since we apply BALANCE only when $\Delta_{L}\geq 2\delta_{L}$, starting with $\Delta_{L}\in\{\delta_{L},\delta_{L}+1\}$, there are at least $\delta_{L}N/4=\Omega(m_{t}+n)=\Omega(m+n)$ edge insertions before that.
  • $\tt{RECOMPUTE}$. The cost of RECOMPUTE is $O(n+m)$. Consider the potential function $|m-m_{t}|$. After a call to RECOMPUTE it is always 00. Now, if the call to RECOMPUTE followed an edge insertion, then $m=2m_{t}+n$. If the call to RECOMPUTE followed an edge deletion, then $m=0.5m_{t}\geq n$. In both cases, the difference between the potentials after and before the call is negative and proportional to the cost. Therefore, the amortized cost is $O(1)$.
5.3 Blowup-conductance tradeoff
Abboud and Wallheimer [3] observed that in some cases, e.g., for exponential-time problems and approximation problems, even a linear blowup might be too costly. Hence, they suggested a variant of their core gadget that gives a tradeoff between the number of vertices and the conductance of $G_{exp}$. Their idea can be adapted to our core gadget as follows. Given $0<\varepsilon\leq\delta\leq 1$, modify the core gadget by setting $N=\lceil\delta n(1+o(1))\rceil$, and for every $v\in V$, instead of adding $\deg_{G}(v)+3$ edges from $v$ to $L$, add $\lceil\varepsilon\deg_{G}(v)\rceil+3$ such edges. By Lemma 10, the resulting graph is an $\Omega(\varepsilon)$-expander. Moreover, the blowup in the number of vertices is $2N\leq 2\delta n+O(1)$, and the blowup in the number of edges is $2\varepsilon m+3n$. Rescaling appropriately, we obtain that given any $0<\varepsilon\leq\delta\leq 1$, we can construct an $\Omega(\varepsilon)$-expander with a blowup of $\delta n$ in the number of vertices and $\varepsilon m+3n$ in the number of edges.
5.4 Bipartiteness-preserving core gadget
In some instances, we would like our core gadget to preserve bipartiteness, e.g., when the problem is defined for bipartite graphs. Note that our core gadget does not preserve bipartiteness because the endpoints of any odd-length path in $G$ may share a common neighbor in $L$, thus closing an odd cycle and making $G_{exp}$ non-bipartite. Therefore, we suggest a variant of the core gadget that, given a bipartite graph $G$, outputs a bipartite, $\Omega(1)$-expander $G_{bexp}:=(V_{bexp},E_{bexp})$.
Given a bipartite graph $G=(A\cup B,E)$, where $|A|=|B|=n$, let us modify the core gadget as follows. Instead of adding edges from $A\cup B$ to $L$, add edges from $A$ to $L$ and from $B$ to $R$. Namely, for every $v\in A$, add $\deg_{G}(v)+3$ edges to $L$, and for every $v\in B$, add $\deg_{G}(v)+3$ edges to $R$. The rest of the construction stays the same. Namely, we construct a $d$-regular expander $X$ between $L$ and $R$ for the same $d$ as before. Clearly, the blowup in the number of edges and vertices is still linear, and the graph is bipartite, with sides $A\cup R$ and $B\cup L$. In addition:
Claim 13.
$G_{bexp}$ is a bipartite $\Omega(1)$-expander.
To see why this claim holds, let $A\cup B$ play the role of $V$, and let $L\cup R$ play the role of $L$ in the original proof of Lemma 9.
6 Direct-WTERs for Max-Cut, Densest Subgraph, and Graphical OMv
In this section, we begin by further developing this line of research by providing Direct-WTERs for additional problems not considered in [3]. Let us begin with the Max-Cut problem.
6.1 Direct-WTER for Max-Cut
In the Max-Cut problem, the goal is to compute the maximum cut in a graph $G$, which we denote by $MC(G):=\max_{S\subseteq V}e(S,V\setminus S)$. To make the graph an $\Omega(1)$-expander, simply applying the core gadget does not work because it might affect the maximum cut in $G_{exp}$ unpredictably. To this end, we add a gadget that ensures that any maximum cut in $G_{exp}$ will separate $L$ from $R$. In addition, we modify the core gadget so that every vertex in $G$ will have the same number of neighbors in $L$ and $R$. These two gadgets together ensure that the vertices in $G$ do not get a preference to be on either side of the cut due to symmetry. Consequently, the maximum cut in $G_{exp}$ induces a maximum cut in $G$.
Let us now describe the full reduction. Given $0<\varepsilon\leq 1$, we apply the core gadget with a tradeoff between the conductance and the blowup, as described in Appendix 5.3. Specifically, we augment $G$ to an $\Omega(\varepsilon^{2})$-expander, with a blowup of $2N=\varepsilon n$ in the number of vertices. This is achieved (roughly) by picking $N=\varepsilon/2n$, and adding for every $v\in V$, $\varepsilon^{2}\deg_{G}(v)+3$ neighbors in $L$. To make the construction more symmetric (so it does not affect the maximum cut), we will do the same thing with respect to $R$, i.e., add for $v\in V$ the same number of neighbors as in $L$.
Observe that for every vertex in $L\cup R$, the number of neighbors it has in $V$ is at most $d\leq\frac{2\varepsilon^{2}m+4n}{N}=4\varepsilon m/n+O(1)\leq 4\varepsilon n+O(1)$. In addition, $d$ is also the degree of the expander between $L$ and $R$. Now we add two bi-cliques of size $N\times 3d$ as follows. Add two sets of $F_{L}$ and $F_{R}$ containing $3d$ vertices each, and all $L$-to-$F_{L}$ and $R$-to-$F_{R}$ edges. The purpose of this gadget is to ensure that $L$ and $R$ are separated in a maximum cut. Denote the resulting graph by $G_{exp}$ (see Figure 3).
Figure 3: Our Direct-WTER for the Max-Cut problem.
Observe that the blowup in the number of vertices is $2N+6d\leq 25\varepsilon n+O(1)$. In addition, we claim that the additional gadgets we add to the core gadget, namely the $V$-to-$R$ edges and the bi-cliques, do not ruin the graph’s expansion. In particular, $G_{exp}$ is an $\Omega(\varepsilon^{2})$-expander. To see why, observe that the induced graph on $L\cup F_{R}\cup F_{L}\cup R$ is a bipartite, $\Omega(d)$-edge expander, hence the proof of Lemma 10 holds for it as well (by letting $L\cup R$ play the role of $L$). The next claim shows that the maximum cut in $G_{exp}$ encodes the maximum cut in $G$.
Claim 14.
The maximum cut in $G_{exp}$ is $MC(G_{exp})=MC(G)+7dN$.
Proof.
Let $(S,V_{exp}\setminus S)$ be a maximum cut in $G_{exp}$. We call the vertices of $S$ red and the vertices of $V_{exp}\setminus S$ blue. Consider the following observations:
  1. 1.
    $F_{L}$ must be monochromatic because otherwise, we could improve the cut by giving $F_{L}$ the opposite color of the majority color in $L$. A similar observation applies to $F_{R}$ as well.
  2. 2.
    $L$ is monochromatic and has the opposite color to $F_{L}$. To see why, assume w.l.o.g. that $F_{L}$ is red, and for contradiction, assume that some $x\in L$ is also red. By changing the color of $x$ to blue, we gain $|F_{L}|=3d$ edges and lose at most $\deg_{V}(x)+\deg_{R}(x)\leq 2d$ edges, a contradiction to $S$ being a maximum cut. A similar observation applies to $F_{R}$ as well.
  3. 3.
    $L$ and $R$ have different colors. To see why, assume w.l.o.g. that $L$ is red, and assume for contradiction that $R$ is also red. Based on the previous observation, $F_{L}$ is blue. By flipping the colors of $L$ and $F_{L}$, we gain the $L$-to-$R$ edges; there are $N\cdot d$ such edges, and lose the $L$-to-$V$ edges; there are at most $N\cdot d$ such edges. Hence, we can assume without loss of generality that this observation holds.
Now, we claim that $S\cap V$ is a maximum cut in $G$. To see why, observe that since $L$ and $R$ have different colors, and for every $v\in V$, we have $\deg_{L}(v)=\deg_{R}(v)$, then replacing $S\cap V$ with any other cut in $V$ will only change the number of edges cut inside $G$. Therefore, we get $MC(G_{exp})=MC(G)+N(|F_{L}|+|F_{R}|+d)=MC(G)+7dN$. ∎
6.2 DD-WTER for Densest Subgraph
In the Densest Subgraph problem, we define the density of non-empty set $S\subseteq V$ to be $\rho(S):=m_{S}/|S|$, where $m_{S}$ is the number of edges in the subgraph induced by $S$. The goal is to compute $\rho(G):=\max_{S\subseteq V}\rho(S)$. In our DD-WTER, we will make use of the following claim.
Claim 15.
Any set $S^{*}$ that maximizes $\rho$ in a graph with $m$ edges and $n$ vertices does not contain any vertex of degree less than $m/n$.
Proof.
Assume for contradiction that $S^{*}$ contains a vertex $u$ of degree $\deg(u)<m/n$. Define $S=S^{*}\setminus\{u\}$. Then, since $m/n=\rho(V)\leq\rho(S^{*})=m_{S^{*}}/|S^{*}|$, we get:
$\rho(S)\geq\frac{m_{S^{*}}-\deg(u)}{|S^{*}|-1}>\frac{m_{S^{*}}-\frac{m}{n}}{|S^{*}|-1}\geq\frac{m_{S^{*}}-\frac{m_{S^{*}}}{|S^{*}|}}{|S^{*}|-1}=\rho(S^{*}),$
which contradicts $S^{*}$ being a maximizer. ∎
In addition, we assume that the density in $G$ is sufficiently large, namely, $m>42n$ [^13] For clarity, we do not attempt to optimize this constant, although it can be reduced further. To make the graph an $\Omega(1)$-expander, we apply the core gadget in a way that only introduces vertices of degree smaller than $m_{exp}/n_{exp}$, where $m_{exp}$ and $n_{exp}$ are the number of edges and vertices in $G_{exp}$, respectively. To this end, consider the graph obtained by applying the core gadget with a tradeoff between the conductance and the blowup, as described in Appendix 5.3. Specifically, we pick the parameters such that in $G_{exp}$, every $v\in V$ is connected to $\lceil\varepsilon\deg_{G}(v)\rceil+3$ vertices in $L$, and $N=|L|=|R|=n(1+o(1))$. The conductance of this graph is $\Omega(\varepsilon)$. We pick $\varepsilon$ to be some sufficiently small constant, which will be determined later.
Claim 16.
The maximum density in $G_{exp}$ is $\rho(G_{exp})=\rho(G)$.
Proof.
We will show that $m_{exp}/n_{exp}>\mu$, where $\mu$ is the maximum degree in $L$ (and therefore also in $L\cup R$). Hence, by Claim 15, it will follow that any maximum density subgraph in $G_{exp}$ does not contain vertices from $L\cup R$, so it must induce a maximum density subgraph in $G$.
Note that the number of $V$-to-$L$ edges is bounded by $2\varepsilon m+3n\leq e(V,L)\leq 2\varepsilon m+4n$, and that the degree of the expander between $L$ and $R$ is $d=\lceil e(V,L)/N\rceil$. Hence, we have $m_{exp}=m+e(V,L)+Nd\geq m+(2\varepsilon m+3n)+(2\varepsilon m+3n)=(1+4\varepsilon)m+6n$. In addition, note that the maximum degree in $L$ is:
$\mu\leq 2d\leq 2\left(\frac{e(V,L)}{N}+1\right)\leq\frac{4\varepsilon m+9n+o(n)}{n}.$
Now, observe that the inequality:
$\mu\leq\frac{4\varepsilon m+9n+o(n)}{n}<\frac{(1+4\varepsilon)m+6n}{3n+o(n)}\leq\frac{m_{exp}}{n_{exp}},$
holds when $n$ is sufficiently large, $01-8\varepsilon>$ 0, and $21n/(1-8\varepsilon)<m$. For example, by picking $\varepsilon<1/16$, we get that $\mu<m_{exp}/n_{exp}$ when $m>42n$. ∎
To adapt the Direct-WTER to a DD-WTER, we replace the core gadget with the dynamic core gadget. However, note that in the dynamic core gadget, there are some variations in the degrees due to lazy updates and rebalances. Nonetheless, these variations can be compensated for by picking a smaller $\varepsilon$. To see why, note that in the dynamic graph, we always have $\varepsilon m+3n\leq e(V,L)\leq 4\varepsilon m+4n$. Hence, $m_{exp}\geq(1+2\varepsilon)m+6n$ and $\mu\leq(4\varepsilon m+9n+o(n))/n$. This implies that the above inequality holds when $n$ is sufficiently large, $01-22\varepsilon>$ 0, and $21n/(1-22\varepsilon)<m$. By picking $\varepsilon=1/44$ we obtain the result.
6.3 DD-WTER for Graphical OMv instances
In this subsection, we demonstrate how our core gadget from Section 4 can be used to prove OMv-hardness to various OMv-hard problems by making the typical “OMv-hard” instances of a problem $\Omega(1)$-expanders. The definitions of the OMv and OuMv problems and the OMv Conjecture appear in Appendix D.
A Graphical OMv instance is constructed from an OuMv instance as follows. Given a $k\times k$ binary matrix $M$, construct bipartite graph $G_{M}:=(A\cup B,E)$, where $A$ and $B$ are equally-sized parts, denoted by $A=\{a_{1},\ldots,a_{k}\}$ and $B=\{b_{1},\ldots,b_{k}\}$. The edges of $G_{M}$ are defined according to the $1$’s of the matrix, i.e., $E:=\{a_{i}b_{j}|M[i,j]=1\}$. Next, add some problem-specific gadgets to $G_{M}$: for many OMv-hard problems, such as $st$-SubConn, $st$-SP, and more, the gadgets consist of $O(k)$ vertices and $O(k)$ edges that are connected to $L\cup B$ in a certain (dynamic) way.
Making Graphical OMv instances $\Omega(1)$-expanders.
To make such instances $\Omega(1)$-expanders, we apply the bipartiteness-preserving core-gadget on $G_{M}$ before adding the problem-specific gadgets. In some cases, as we will soon demonstrate, such gadgets preserve expansion, or they can be easily adapted to preserve expansion. Therefore, this adaptation proves that their OMv-hard instances are $\Omega(1)$-expanders. Let us demonstrate this technique for the problem of $st$-SP (see Appendix D for its definition).
Proof of Proposition 5.
We will prove the proposition for an easier problem called $st$-SP ($3$ vs. $5$), where the goal is only to distinguish between $dist(s,t)=3$ and $dist(s,t)\geq 5$. Henzinger et al. [24] proved a lower bound to this problem via a reduction from OuMv: construct $G_{M}$, add vertices $s$ and $t$ to $G_{M}$, and then update the edges between $s$-to-$A$ and $t$-to-$B$ according to the input vectors. It then followed that whenever $u^{T}Mv=1$, then $dist(s,t)=3$, and whenever $u^{T}Mv=0$, $dist(s,t)\geq 5$. By picking $k=\sqrt{m}$, the graph has $O(m)$ edges, and the lower bound the follows was $m^{1/2-\varepsilon}$ per update and $m^{1-\varepsilon}$ per query. We now modify their construction as follows.
Given $M$, apply the bipartiteness-preserving core gadget to $G_{M}$ before adding vertices $s$ and $t$. Then, pick a non-edge in the expander $X$, i.e., $xy\notin E(X)$ for some $x\in L$ and $y\in R$ arbitrarily. The purpose of this modification is to ensure that $s$ and $t$ are connected to the graph without introducing a path of length $<5$ between them. Now, we procceed as in [24]; namely, given vectors $u=(u_{1},u_{2},\ldots,u_{k})$ and $v=(v_{1},v_{2},\ldots,v_{k})$, we update the graph by adding the edges $sa_{i}$ iff $u_{i}=1$, and $tb_{j}$ iff $v_{j}=1$. Note that $u^{T}Mv=1$ iff $dist(s,t)=3$, and otherwise $dist(s,t)\geq 5$. In addition, we claim that the graph is an $\Omega(1)$-expander at all times. This follows from the next claim.
Claim 17.
If $G$ is a $\phi$-expander for some $\phi>0$, then adding a vertex $v$ to $G$ and connecting it to $\ell\geq 1$ arbitrary vertices, results in a graph $H$ which is a $\phi/4$-expander.
In our setting, we add two vertices to $G_{M}$, and at all times, each of them is connected to at least one vertex in $G_{M}$. Therefore, since $G_{M}$ is a $\phi$-expander for some constant $\phi$, then the resulting graph is a $\phi/16$-expander at all times according to this claim.
Proof of Claim 17.
Consider a non-empty cut $S\subseteq V(H)$. Denote by $S_{V}=S\cap V$, and assume without loss of generality that $vol_{G}(S_{V})\leq vol_{G}(V\setminus S_{V})$ (we can assume so because it holds either for $S$ or its complement). This assumption is useful since we then have $e_{G}(S_{V},V\setminus S_{V})\geq\phi vol_{G}(S_{V})$, as $G$ is a $\phi$-expander. Now, we split into the following cases. If $S$ does not contain $v$, i.e. $S_{V}=S$, then $vol(S)\leq vol_{G}(S)+|S|\leq 2vol_{G}(S)$, therefore $e(S,V(H)\setminus S)\geq e_{G}(S,V\setminus S)\geq\phi/2\cdot vol(S)$. If $S$ contains $v$, then $vol(S)\leq\ell+2vol_{G}(S_{V})$. Now, if $\ell\leq 2|S_{V}|$, then $vol(S)\leq 4vol_{G}(S_{V})$, and as in the previous case we have $e(S,V(H)\setminus S)\geq e_{G}(S_{V},V\setminus S_{V})\geq\phi/4\cdot vol(S)$. If $\ell>2|S_{V}|$, then $e(\{v\},V\setminus S_{V})\geq\ell/2$. Therefore, we get
$e(S,V(H)\setminus S)\geq\ell/2+\phi vol_{G}(S_{V})\geq\phi/2(\ell+vol_{G}(S_{V}))\geq\phi/4\cdot vol(S)$
Hence, assuming the OMv Conjecture, there is no algorithm for $st$-SP ($3$ vs. $5$) on $\Omega(1)$-expanders, whose preprocessing time is polynomial, update time $m^{1/2-\varepsilon}$, and query time $m^{1-\varepsilon}$, for any $\varepsilon>0$. ∎
7 Derandomized Direct-WTERs for Additional Problems
In this section, we show how the deterministic core gadget and its dynamic adaptation can be combined with additional gadgets from [3] to obtain deterministic Direct-WTERs and DD-WTERs to all problems that were considered in [3], and for some other related problems.
Proposition 18.
The following problems admit deterministic Direct-WTERs: Maximum Matching, Minimum Vertex Cover, $k$-Clique Detection, $k$-Clique Counting, Max-Clique, Minimum Dominating Set, and $H$-Subgraph Detection ($m=O(m)$ and $H$ does not contain pendant vertices).
The following problems admit DD-WTERs: Maximum Matching, Bipartite Perfect Matching, $k$-Clique Detection, $k$-Clique Counting, and $H$-Subgraph Detection ($m=O(m)$ and $H$ does not contain pendant vertices).
We remark that for the $k$-Clique Detection and $k$-Clique Counting problems, our additional gadget differs from the one in [3]. Hence, we explain them in more detail.
Maximum Matching, Bipartite Perfect Matching, Minimum Vertex Cover.
The Direct-WTER for Maximum Matching and Minimum Vertex Cover are as follows. Given $G$, apply the core gadget to obtain $G_{exp}$. Then, for every $w\in L\cup R$, add a pendant vertex connected to $w$ (that is, a degree-$1$ vertex $w^{\prime}$ whose only neighbor is $w$). Denote by $MM(\cdot)$ the cardinality of the maximum matching in a graph. By [3, Fact 16], the cardinality of the maximum matching in the resulting graph is $MM(G)+2N$. A similar claim holds for minimum vertex cover. By [3, Lemma 17], the graph is an $\Omega(1)$-expander. Notice for the Minimum Vertex Cover problem, which is an exponential time problem, we will need to use the blowup-conductance tradeoff, resulting in an $\Omega(\varepsilon)$-expander with $n(1+\varepsilon+o(1))$ vertices.
For the problem of Bipartite Perfect Matching, both the input and output of the reduction are bipartite graphs. Hence, we employ the bipartiteness-preserving core gadget on $G$ to obtain $G_{exp}$ and then add pendant vertices as before. Note that the resulting graph is bipartite and contains a perfect matching if and only if $G$ contains a perfect matching.
The adaptation of this reduction to the dynamic setting for Maximum Matching is straightforward, as we need to replace the core gadget with the dynamic core gadget. For Bipartite Perfect Matching, we combine the bipartiteness-preserving variant of the core gadget to the dynamic setting by replacing the role of $L$ with $L\cup R$. Hence, Bipartite Perfect Matching admits a DD-WTER.
$k$-Clique Detection, $k$-Clique Counting
In the following reduction, we assume w.l.o.g. that the graph contains no isolated vertices, as isolated vertices can be easily dealt with. Begin by taking a copy of $V$, denoted $V_{ind}=\{v_{ind}\mid v\in V\}$, in which all vertices are independent. In addition, add the following set of edges $\{uv_{ind}|uv\in E\}$. Next, apply the core gadget on the graph, but apply it only to $V_{ind}$ instead of $V\cup V_{ind}$. Apply the core gadget on a graph whose vertex set is $V_{ind}$ $V$ and add edges only from $V_{ind}$ to $L$. Denote the resulting graph by $\bar{G}:=(\bar{V},\bar{E})$. See Figure 4. Observe that for every $k\geq 3$, all $k$-cliques in $G$ appear in $\bar{G}$. In addition, any $k$-clique in $\bar{G}$ that is not contained in $G$ must contain $k-1$ vertices in $V$, and a single vertex $w_{ind}\in V_{ind}$. In this case, there is a corresponding $k$-clique in $G$, obtained by replacing $w_{ind}$ with $w$. From this observation, we get the following corollary:
Corollary 19.
Let $c_{k}$ denote the number of $k$-cliques in $G$. Then the number of $k$-cliques in $\bar{G}$ is $(k+1)c_{k}$.
Therefore, this reduction is applicable for $k$-Clique Detection and $k$-Clique Counting.
Let us now prove that this additional gadget of constructing $V_{ind}$ before applying the core gadget does not ruin the expansion by much.
Claim 20.
$\bar{G}$ is an $\Omega(1)$-expander.
The proof of this claim follows along the lines of the proof of Lemma 9. In particular, we split into two cases as follows. If a large volume is concentrated in $V$, then there must be many outer edges from $V$ to $V_{ind}$. We then employ Lemma 9 to argue that there are many outer edges overall. If there is not much volume in $V$, the Lemma 9 suffice to prove the claim.
Proof.
Consider any cut $S=S_{V}\cup S_{ind}\cup S_{L}\cup S_{R}$, where $S_{V}\subseteq V,S_{ind}\subseteq V_{ind},S_{L}\subseteq L,S_{R}\subseteq R$, such that $|S_{L}\cup S_{R}|\leq N$. Since there are no isolated vertices, then $vol(S)>0$. Observe that by Lemma 9, we have
$e(S_{V},\bar{V}\setminus S_{V})\geq e(S_{ind},L\setminus S_{L})+e(S_{L},R\setminus S_{R})\geq\phi vol(S_{ind}\cup S_{L}\cup S_{R}),$
for some constant $\phi>0$. We now split into two cases as follows.
  • $vol(S_{V})\leq 4vol(S_{ind})$. In this case, we have $vol(S)\leq 5vol(S_{ind}\cup S_{L}\cup S_{R})$. Hence, by the above observation, we have $e(S,\bar{V}\setminus S)\geq\phi vol(S)/5$, and we are done.
  • $vol(S_{V})>4vol(S_{ind})$. In this case, we have:
    $e(S_{V},V_{ind}\setminus S_{ind})=e(S_{V},V_{ind})-e(S_{V},S_{ind})>vol(S_{V})/2-vol(S_{V})/4=vol(S_{V})/4.$
    Hence, we get $e(S,\bar{V}\setminus S)\geq vol(S_{V})/4+\phi vol(S_{ind}\cup S_{L}\cup S_{R})\geq\phi/4vol(S)$, and we are done.
Adapting this reduction to the dynamic setting is by replacing the core gadget with the dynamic one, and making two edge updates to $uv_{ind},u_{ind}v$ for every edge update $uv$ in $G$. We remark that to deal with isolated vertices in the dynamic setting, we can ignore deletions that remove the last edge incident to a vertex. Such edges can not participate in a $k$-Clique, hence the correctness is preserved.
Figure 4: Our Direct-WTER for $k$-Clique Detection, $k$-Clique Counting, and Max Clique. The dashed edges correspond to edges in $G$.
$H$-Subgraph Detection ($m=O(n)$ and $H$ does not contain pendant vertices)
Apply the core gadget to obtain a graph $G_{exp}$. Then, subdivide each edge in $E(V,L)$ and $E(L,R)$ into a path of length $|V(H)|/2$. By [3, Claim 21], any copy of $H$ in the resulting graph must be a copy of $H$ in $G$. By [3, Claim 22], if $|V(H)|=O(1)$ then the resulting graph is an $\Omega(1)$-expander. As observed in [3], the blowup in the number of vertices is $\Omega(m)$. Hence, it is only applicable to graphs with $m=O(n)$ edges. Adapting this reduction to the dynamic setting is done by replacing each edge insertion or deletion in $E(V,L)\cup E(L,R)$ by insertion of a path, resulting in a slowdown of $O(|V(H)|)=O(1)$.
Max-Clique and Minimum Dominating Set
The gadget employed for these problems is similar to the one we used for $k$-Clique Detection. Namely, we will construct the set $V_{ind}$ and then apply the core gadget. However, since we now deal with exponential problems, there is a need to reduce the size of $V_{ind}$ to $\varepsilon n$. In [3, Section 4.6], the authors show a randomized algorithm for constructing a subset of vertices $Q\subseteq V$ of size $\varepsilon n$, that “hits” all sparse cuts in $G$. Namely, for any cut $S\subseteq V$, either $vol_{G}(S\cap Q)\geq f(\varepsilon)\cdot vol(S)$, or $e_{G}(S,V\setminus S)>f(\varepsilon)vol(S)$, where $f(\varepsilon)=\frac{\varepsilon^{2}}{10\log(1/\varepsilon)}$. The only randomized component of this construction is to pick a random subset $Q\subseteq V$ of cardinality $\varepsilon n$, such that for every vertex $v\in V$ for which $\deg_{G}(v)>\frac{\log(1/\varepsilon)}{\varepsilon}$, $Q$ hits an $\varepsilon$-fraction of the neighbors of $v$.This component can be derandomized, e.g., by exhaustive search, incurring a (affordable) slowdown of $O(2^{\varepsilon n})$ time.
For the problem of Max-Clique, the reduction is by constructing $Q_{ind}=\{u_{ind}|u\in Q\}$ and $E(V,Q_{ind})=\{vu_{ind}|u\in Q\}$, and then applying the generalized core gadget with respect to $Q_{ind}$ to obtain a conductance-blowup tradeoff. Namelt, we connect every $w_{ind}\in Q_{ind}$ to $\lceil\varepsilon\deg(w_{ind})\rceil+3$ neighbors in $L$, and set $N=\lceil\varepsilon n\rceil+o(n)$. Denote the resulting graph by $\bar{G}$. It follows from our previous observations that any $k$-clique in $\bar{G}$ corresponds to a $k$-clique in $G$. Hence, the maximum clique in $\bar{G}$ has the same size as the maximum clique in $G$. The blowup in the number of vertices in $\bar{G}$ is $2\varepsilon n+o(n)$. In addition:
Claim 21.
$\bar{G}$ is an $\Omega(f(\varepsilon))$-expander.
The proof splits into the following cases. Either the cut $S_{V}$, is hit by $Q$, in which case there are many edges to $Q_{ind}$, and the proof continues by combining the proofs of Claim 20 and Lemma 10. Otherwise, by the properties of $Q$, there must be many outer edges inside $V$ that account for $vol(S_{V})$. Then, the proof follows from the previous results.
Proof.
Consider any cut $S=S_{V}\cup S_{Q}\cup S_{L}\cup S_{R}$, where $S_{V}\subseteq V,S_{Q}\subseteq Q_{ind},S_{L}\subseteq L,S_{R}\subseteq R$, such that $|S_{L}\cup S_{R}|\leq N$. Clearly, $vol(S)>0$, assuming there are no isolated vertices in $G$. Observe that by Lemma 10, we have that
$e(S_{Q}\cup S_{L}\cup S_{R})\geq\varepsilon\phi vol(S_{Q}\cup S_{L}\cup S_{R})$
for some constant $\phi>0$. We split into the following cases:
  • $vol_{G}(S_{V}\cap Q)<f(\varepsilon)vol_{G}(S_{V})$. In this case, by the properties of $Q$ we have that $e(S_{V},V\setminus S_{V})\geq f(\varepsilon)vol_{G}(S_{V})\geq f(\varepsilon)vol(S_{V})/2$. Hence,
    $\begin{split}&e(S,\bar{V}\setminus S)\geq f(\varepsilon)vol(S_{V})/2+e(S_{Q},L\setminus S_{L})+e(S_{L},R\setminus S_{R})\geq\\ &f(\varepsilon)vol(S_{V})/2+\phi vol(S_{Q}\cup S_{L}\cup S_{R})\geq\phi f(\varepsilon)vol(S)/2.\end{split}$
  • $vol_{G}(S_{V}\cap Q)\geq f(\varepsilon)vol_{G}(S_{V})$. In this case, by the construction we have that $e(S_{V},Q)\geq f(\varepsilon)vol_{G}(S_{V})\geq f(\varepsilon)vol(S_{V})/2$. We split into two more cases, as follows. If $vol(S_{V})<4vol(S_{Q})/f(\varepsilon)$, then by the above observation we have:
    $e(S_{Q}\cup S_{L}\cup S_{R})\geq\phi vol(S_{Q}\cup S_{L}\cup S_{R})>\phi f(\varepsilon)vol(S)/4.$
    Otherwise, we have that $e(S_{V},Q\setminus S_{Q})$ must be large:
    $e(S_{V},Q\setminus S_{Q})=e(S_{V},Q)-e(S_{V},S_{Q})\geq f(\varepsilon)vol(S_{V})/2-f(\varepsilon)vol(S_{V})/4\geq f(\varepsilon)vol(S_{V})/4.$
    As in the previous case, we get $e(S,\bar{V}\setminus S)\geq f(\varepsilon)\phi vol(S_{V})/4$.
The reduction for Minimum Dominating Set is the same, but with the additional gadget of adding pendant vertices to the vertices in $L$. Denote the cardinality of the Minimum Dominating Set by $MDS(\cdot)$. By [3, Fact 23], it follows that the cardinality of the minimum dominating set in $\bar{G}$ is $MDS(G)+|L|=MDS(G)+N$. By [3, Lemma 17], the graph is an $\Omega(\varepsilon)$ expander.
8 Distributed Core Gadget
In the distributed CONGEST model, the input graph plays the additional role of serving as the communication network over which nodes can send messages of length $O(\log n)$. Expander decompositions in CONGEST have been utilized to obtain major breakthroughs, e.g., for Triangle Enumeration [12]. We remark that the current techniques for Direct-WTERs cannot be used in the CONGEST model. The reason is that those techniques augment the graph with additional edges, but in CONGEST, the edges are also communication links that cannot be added (they could be simulated, but that would incur large congestion). This is not so surprising since, in this model, expanders tend to be provably easier than worst-case graphs. For example, constructing a spanning tree in $n^{o(1)}$ CONGEST rounds [19], while there is a $\Omega(\sqrt{n}/\log n)$ lower bound for general graphs [35].
We focus on other related distributed models that may have been less studied in this context but may also benefit from expander decompositions. One such model is the CONGESTED-CLIQUE model, in which the input graph is separated from the communication network, which is a clique. Namely, in every round, each node can send $O(\log n)$ bits to every other node. Another related model which we consider is the MPC (Massively Parallel Computation) model. In MPC, the input graph is distributed across $M$ machines arbitrarily, each machine has a local space of $S$ words (where a word has a size of $O(\log n)$ bits), and we assume that the total amount of space is $M\cdot S=\Theta(n+m)$. In every MPC round, each machine can send a message consisting of $S$ words to another machine and reads at most $S$ words sent by other machines. We focus on sublinear-MPC, where $S=n^{\delta}$ for some constant $10<\delta<$ 1.
We initiate the study of WTERs in these two models. Our results imply that the core gadget can be adapted to both models, resulting in Direct-WTERs in CONGESTED-CLIQUE and Sublinear-MPC for all problems that were considered in the static and dynamic settings. Namely:
Proposition 22.
The following problems admit a deterministic Direct-WTER in CONGESTED-CLIQUE and Sublinear-MPC: Maximum Matching, $k$-Clique Detection, $k$-Clique Counting, Densest Subgraph ($m\geq 48n$), Minimum Dominating Set, Minimum Vertex Cover, Max-Cut, $H$-Subgraph Detection (where $m=\tilde{O}(n)$ and $H$ does not contain pendant vertices).
8.1 A core gadget in CONGESTED-CLIQUE and Sublinear-MPC
CONGESTED-CLIQUE
We will refer to the basic components of the input graph as vertices and to the basic components of the communication network as nodes. At the beginning, every node corresponds to a vertex. The goal is to simulate the graph $G_{exp}$. Some nodes in the network will play two roles: the role of their original vertex (some $v\in V$) and the role of some vertex in $L\cup R$. In the first round, a special node is chosen to receive all the degrees of all the vertices in the graph. The special node simulates Round-Robin locally and computes indices $\{i_{v}\}_{v\in V}$ corresponding to the starting positions given by the Round-Robin for every vertex in $V$ (i.e. $i_{v}$ is the position such that vertex $v$ is connected in $L$ to $x_{i_{v}},\ldots,x_{i_{v}+\deg_{G}(v)+2}$). It then sends them back in the next round so that every node corresponding to $v\in V$ has a list of neighbors that $v$ has in $G_{exp}$. In addition, the special node computes the number of edges in $G$ and sends them to the nodes playing the role of $L\cup R$. Each of these nodes can compute the expander $X$ locally. In the final round, each node corresponding to $v\in V$ sends its ID to the nodes corresponding to its neighbors in $L$. The number of rounds this algorithm needs is $O(1)$, and the result of this algorithm is that every node that corresponds to $v\in V_{exp}$ knows the edges incident at $v$ in $G_{exp}$.
Sublinear-MPC
The goal is to compute the edges of $G_{exp}$ in $O(1)$ rounds so that, in the end, they are distributed across the machines. We will assume that the total space in the system suffices to store $G_{exp}$, which has a linear size in $G$. If we try to adapt the above algorithm to the MPC model, the first difficulty encountered is that we can not designate a special machine to receive all the degrees and simulate Round-Robin. That is because each machine is limited to receive at most $n^{\delta}$ words in every round. Hence, we execute a “parallel Round-Robin” as follows. We will assume that each vertex has a corresponding machine, which knows its degree in $G$ (multiple vertices can correspond to the same machine, which stores their degrees while respecting the space constraint). This can be computed deterministically in $O(1)$ rounds, following from [22].
To parallelize the Round-Robin algorithm, we simulate a communication tree of size $O(n)$, degree $n^{\delta}$, and depth $O(1/\delta)$, whose leaves correspond to $V$, and every node in the tree stores the sum of degrees of all the vertices in its subtree. To this end, the leaves of the tree are simulated using the machines that store the degrees (at most $O(M)$), Moreover, we will designate additional $O(n^{1-\delta})$ machines that will act as internal nodes; hence, the total number of machines is $O(M+n^{1-\delta})=O(M)$. The values (degree sums) of the tree are computed in $O(1/\delta)$ rounds, starting from the leaves bottom-up. In particular, Clearly, the sum at the root is $2m$.
In the next phase, we compute the indices $\{i_{v}\}_{v\in V}$ corresponding to the starting points given by the Round-Robin algorithm. This is done top-bottom: the root is given the degree-sums of its children and simulates Round-Robin on “super-nodes” corresponding to nodes in disjoint subtrees, whose degrees are the degree-sums of the subtrees. It then sends the starting indices to the super-nodes. Each super-node then splits its degree-sum to the sum of degrees of its children and repeats the process, starting the Round-Robin simulation from its starting index computed by its parent. This process terminates when the leaves, corresponding to $V$, receive the set of indices $\{i_{v}\}_{v\in V}$. This takes $O(1/\delta)$ rounds. The proof of correctness that the computed indices indeed correspond to the indices given by the Round-Robin algorithm is by induction on the levels of the tree.
The system now has a set of machines that each store a subset of degrees of vertices of $V$ and their starting indices of neighbors in $L$, such that each vertex in $V$ corresponds to some machine. To compute the edges of $G_{exp}$ explicitly:
  • We designate $O(M)$ machines that receive starting indices and degrees and compute the $V$-to-$L$ edges accordingly. This can be done because there are at most $O(m+n)$ such edges.
  • We designate $O(M)$ machines that compute and store the edges of $X$. This can be done since $X$ has size $O(m+n)$.
The resulting state of the system is that every edge of $G_{exp}$ is stored in some machine.
9 Acknowledgements
We would like to thank Thatchaphol Saranurak for clarifying a typo in the mentioned lower bound for the Densest Subgraph problem [24, Table 5].
Appendix A OMv-hardness for Densest Subgraph in Denser Graphs
In this appendix, we provide a modification to a reduction by Henzinger et al. [24], from OuMv to the dynamic Densest Subgraph problem. The goal of this modification is to increase the number of edges in the output graph to be at least $C\cdot n$ for any given integer $C\geq 1$, thus showing that the Densest Subgraph $(m\geq Cn)$ variant of the problem is OMv-hard. Since our DD-WTER for the Densest Subgraph problem only works for sufficiently dense graphs, namely, graphs with $m>42n$ edges, this implies that the lower bound of [24] applies even on $\Omega(1)$-expanders.
The original proof of [24].
Let $M$ be a $N\times N$ binary matrix given as input for OuMv. In [24], the authors define a bit gadget for every $i,j$, such that if $M[i,j]=1$, then the corresponding bit gadget is a path on $6N$ vertices, and if $M[i,j]=0$, then it is an independent set. In addition, they define $N$ row gadgets and $N$ column gadgets as independent sets on $3$ vertices or triangles. One vertex of every row/column gadget is connected to the endpoints of $N$ corresponding bit gadgets. The row/column gadget dynamically changes between a triangle and an independent set according to the input vectors (as defined in the OuMv problem). Denote this graph by $G$. We remark that the maximum density in $G$ is bounded by $2$.
Observation 23.
The maximum density in $G$ is $\rho(G)\leq 2$.
In particular, $|E(G)|\leq 2|V(G)|$. To see why this observation is true, note that every edge in this graph is incident to a degree $2$-vertex. Therefore, for any $S\subseteq V(G)$, we have $m_{S}/|S|\leq 2x/x=2$, where $x$ is the number of vertices of degree $2$ in $S$.
Let us now describe a self-reduction that given any instance $G$ of Densest Subgraph, such that $\rho(G)<C+1/2$ for some integer $C\geq 1$, it produces a graph $G_{C}$ with $|E(G_{C})|\geq C|V(G_{C})|$ edges. This self-reduction adapts trivially to the dynamic setting, and in particular, it can be applied to the dynamic graph described above (which has maximum density $\rho(G)\leq 2$), to produce OMv-hard instances with $m\geq Cn$ edges, for any constant integer $C\geq 2$, in particular $C>42$.
Our modification.
Given some integer $C\geq 1$, attach a $(2C+1)$-clique to every vertex in $G$. Namely, for every $v\in V(G)$, construct a clique containing $v$ and additional $2C$ vertices. Denote the resulting graph by $G_{C}$. Note that the number of vertices in $G_{C}$ is $n=|V(G)|(2C+1)$, and the number of edges is $m\geq|V(G)|\binom{2C+1}{2}=Cn$. The correctness of this modification follows from the next proposition.
Proposition 24.
Let $G$ be a graph such that $\rho(G)<C+1/2$. Then the graph $G_{C}$, obtained from $G$ by attaching a $(2C+1)$-clique to every vertex of $G$, has maximum density $\rho(G_{C})=\rho(G)/(2C+1)+C$.
Hence, one can easily compute the maximum density of $G$ from the maximum density of $G_{C}$. Therefore, the complexity of dynamic Densest Subgraph $(m\geq Cn)$ is at least the complexity (up to $n^{o(1)}$ factors) of dynamic Densest Subgraph when restricted to graphs with density $<C+1/2$, e.g., the graphs output by the reduction in [24]. In particular, we obtain the next corollary.
Corollary 25.
There is no dynamic algorithm for Densest Subgraph ($m\geq Cn$) whose update time is $n^{1/3-o(1)}$ and query time $n^{2/3-o(1)}$, unless the OMv Conjecture fails.
Proof of Proposition 24.
Let $S\subseteq V(G_{C})$ be a set which maximizes $\rho$ in $G_{C}$, i.e. $\rho(S)=\rho(G_{C})$. We can assume w.l.o.g. that $S$ induces a connected component because all connected components induced by $S$ must have the same density. In addition, we have $\rho(S)\geq C$, since any $(2C+1)$-clique has density $C$. The heart of the proof is the following observation: every $(2C+1)$-clique intersected by $S$ must be fully contained in $S$. Essentially, we begin by proving a general claim that holds for all graphs: If some set of vertices intersects a clique but does not contain it, then, we can increase its density by either adding or removing a single vertex from the clique to the set, provided that the intersection occurs at two vertices or more. This implies that any $(2C+1)$-clique intersected by $S$ has either one or $12C+$ 1 vertices in $S$, as otherwise, $S$ would not be a maximizer of $\rho$. Then we use the fact that $\rho(G)<C+1/2$ to show that the former is not possible, i.e., any $(2C+1)$-clique intersected by $S$ is contained in $S$. This shows that $S$ has the form of a densest subgraph in $G$ and attached $(2C+1)$-cliques to every vertex in contains in $G$, yielding the claimed density.
Formally, let us denote by $11\leq k\leq 2C+$ 1 the size of the intersection between $S$ and a $(2C+1)$-clique which is attached to some vertex $v\in V(G)$. We want to show first that $k=1$ or $k=2C+1$. Assume for the sake of contradiction that $11<k<2C+$ 1. Note that $S$ must contain $v$, as otherwise it would not induce a connected component and have density $\geq C$.
Since $1<k$, there must exist some vertex $u\neq v$ in the intersection of $S$ and the attached clique of $v$. Moreover, since $S$ is a maximizer of $\rho$, then the subset $S^{\prime}=S\setminus\{u\}$ has density less than or equal to $\rho(S)$. Hence, we have:
$\rho(S^{\prime})=\frac{m_{S}-(k-1)}{|S|-1}\leq\frac{m_{S}}{|S|},$
which is equivalent to $m_{S}\leq(k-1)|S|$. On the other hand, since $k<2C+1$, there must be some vertex $w\notin S$ which belongs to the attached clique of $v$. Consider the set $S^{\prime\prime}=S\cup\{w\}$. Its density is also less than or equal to $\rho(S)$, hence we have:
$\rho(S^{\prime\prime})=\frac{m_{S}+k}{|S|+1}\leq\frac{m_{S}}{|S|},$
which is equivalent to $m_{S}\geq k|S|$, a contradiction to the previous inequality. Thus, it must be the case that $k=1$ or $k=2C+1$.
Let us now show that $k=2C+1$. For the sake of contradiction, assume that $k=1$, i.e., $S$ contains only $v$. Consider the set $S^{\prime}$, obtained from $S$ by including all the $2C$ remaining vertices of the clique. It has density:
$\rho(S^{\prime})=\frac{m_{S}+\binom{2C+1}{2}}{|S|+2C},$
which is strictly larger than $m_{S}/|S|$, provided that $m_{S}/|S|<C+1/2$. Let us prove then that $m_{S}/|S|<C+1/2$.
Denote by $v_{S}=|S\cap V(G)|$ and by $e_{S}$ the number of edges of $G$ induced by $S$. Then, by the previous claim, we have that $|S|=v_{S}+t\cdot 2C$ and $m_{S}=e_{S}+t\binom{C+1}{2}$, for some $t\geq 0$ representing the number of attached cliques that are fully contained by $S$. Hence, we have that:
$\frac{m_{S}}{|S|}=\frac{e_{S}+t\binom{C+1}{2}}{v_{S}+t\cdot 2C}<C+1/2,$
where the inequality follows from our assumption on $\rho(G)$, namely $e_{S}/v_{S}\leq\rho(G)<C+1/2$. Hence, $\rho(S^{\prime})>\rho(S)$, which contradics $S$ being a maximizer of $\rho$. Therefore $k=2C+1$.
We showed that $S$ is a subset of vertices of $G$ and their corresponding $(2C+1)$-cliques. In particular $|S|=v_{S}(2C+1)$, and $m_{S}=e_{S}+v_{S}\binom{2C+1}{2}$. Therefore:
$\rho(S)=\frac{e_{S}+v_{S}\binom{2C+1}{2}}{v_{S}(2C+1)}=\frac{e_{S}}{v_{S}(2C+1)}+C.$
Note that since $e_{S}/v_{S}\leq\rho(G)$, then $\rho(S)\leq\rho(G)/(2C+1)+C$. However, it must be the case that $v_{S}/e_{S}=\rho(G)$, since this value can be attained by picking any densest subgraph in $G$ and then adding all the attached cliques incident to it. Therefore, $\rho(S)=\rho(G)/(2C+1)+C$. ∎
Appendix B Definition of Dynamic Direct-WTERs
Definition 26 (Dynamic Direct-WTER).
Let $\mathcal{A}$ be a dynamic graph problem, on a dynamic graph $G_{0},G_{1},\ldots$, where $G_{t}:=(V,E_{t})$ is the graph after the $t^{th}$ edge update. Let us denote $n:=|V|$ and $m_{t}:=|E_{t}|$. A Dynamic Direct-WTER to $\mathcal{A}$, is a dynamic algorithm, that: (1) given $G_{0}$, it outputs in $\tilde{O}(n+m_{0})$ time an $\Omega(1)$-expander $G^{exp}_{i_{0}}:=(V_{exp},E^{exp}_{i_{0}})$ with $i_{0}:=0$, and (2) given an edge update to $G_{t-1}$, it computes in amortized $\tilde{O}(1)$ time a sequence of edge updates to $G^{exp}_{i_{t-1}}$, whose length is $i_{t}-i_{t-1}$ for some $i_{t}>i_{t-1}$, resulting in a graph $G^{exp}_{i_{t}}$, such that for every $t\in\mathbb{N}$ we have the following guarantees:
  • The length of the sequence of updates to $G^{exp}_{t-1}$ is amortized $O(1)$, i.e., $i_{t}=O(t)$.
  • Every intermediate graph in $G^{exp}_{i_{t-1}},G^{exp}_{i_{t-1}+1},\ldots,G^{exp}_{i_{t}}$ is an $\Omega(1)$-expander. In addition, the blowup in the number of vertices and edges is $|V^{exp}|\leq K$ and $|E^{exp}_{i_{t-1}+i}|={O}(M)$ for some $K:=L(n,m_{0})$ and $M:=M(n,m_{t})$.
  • The solution $\mathcal{A}(G_{t})$ can be computed in $\tilde{O}(1)$ time from the solution $\mathcal{A}(G^{exp}_{i_{t}})$.
Observe that if problem $\mathcal{A}$ admits a DD-WTER with linear blowup, i.e. $K=O(n)$ and $M=O(N+m_{t})$, then the amortized complexity of the problem on $\Omega(1)$-expanders is equivalent to its amortized complexity on general graphs, up to $poly(\log n)$ factors, and assuming $m=\Omega(n)$. We remark that our definition is w.r.t. the amortized complexity and not worst-case complexity of problems. The reason for this is, that making our Direct-WTERs work both in deterministic and worst-case time per update proved more challenging. Instead, we keep our focus on the deterministic regime, and show Direct-WTERs that work in deterministic amortized $\tilde{O}(1)$ time per update.
Appendix C Explicit Expander Construction
Let us present an explicit, deterministic construction of bipartite $\Omega(1)$-expanders, which builds upon known constructions from the literature. For this part only, we will need to introduce the notion of spectral expanders, which are more common in the literature about explicit constructions.
Graph spectrum.
Consider a $d$-regular, $n$-vertex graph $G$. Denote by $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$ the spectrum of its adjacency matrix. It is known that $\lambda_{1}=d$, and if $G$ is not bipartite then $\lambda_{n}>-d$. There is also a close relationship between $\lambda_{2}$ and the edge-expansion of $G$, which follows by the following theorem.
Theorem 27 (Cheeger’s inequality).
$\frac{d-\lambda_{2}}{2}\leq h_{G}\leq\sqrt{2d(d-\lambda_{2})}$
In particular, if $\lambda_{2}$ is sufficiently small, say $\lambda_{2}\leq 0.1d$, then $G$ is a $\Omega(d)$-edge expander and by Proposition 8 also an $\Omega(1)$-expander.
In our core gadget (Section 4), we will utilize a strongly explicit construction of expanders by Noga Alon [7, Proposition 1.1]. This is a deterministic construction that produces simple, $d$-regular graphs on $n$ vertices, such that $|\lambda_{i}|=O(\sqrt{d})$ for all $i=\in[2,n]$. By strongly explicit, we mean not only the ability to construct the entire graph deterministically, but also the capability to determine the $i^{th}$ neighbor of vertex $u$ in $\tilde{O}(1)$ time, given $i\in[d]$ and $u\in[n]$. In particular, we can construct the entire graph in $\tilde{O}(nd)$ time. For any sufficiently large value of $n$ and $d\geq 3$, this construction is applicable with $n^{\prime}=n(1+o_{n}(1))$ vertices, where the $o_{n}(1)$ term tends to zero as $n$ tends to infinity. For the sake of simplicity, we will assume that $n^{\prime}=n$, as this will not impact our final results.
In fact, we require a strongly explicit construction of a bipartite expander graph since Alon’s graphs are inherently non-bipartite. To achieve this, we adapt the construction through a technique known as a $2$-lift. This adaptation yields a strongly explicit bipartite graph that ensures a bounded second-largest eigenvalue.
Let’s consider a strongly explicit graph $G=(V,E)$, where $|\lambda_{i}|\leq\lambda$ for all eigenvalues $i\in[2,n]$. We establish two copies of $V$, labeled as $V_{1}$ and $V_{2}$. Edges $(u_{1},v_{2})\in V_{1}\times V_{2}$ are introduced if and only if the corresponding edge $(u,v)\in E$. This operation results in a $d$-regular bipartite graph that retains strong explicitness.
According to [10, Lemma 3.1], the eigenvalues of this new bipartite graph correspond to those of $G$ and their additive inverses. Notably, the second largest eigenvalue remains bounded by $\lambda$ through this transformation.
Appendix D Problem Definitions
In this section, we provide the definitions and some background on each problem that appears in the paper.
Maximum Matching and Bipartite Perfect Matching.
In the Maximum Matching problem, the goal is to compute (in the dynamic setting, maintain) the maximum cardinality of any matching, that is, a set of edges $M\subseteq E$ in which no two edges intersect. Relevant background for this problem in the static setting, including background on its expander-case and average-case complexities, can be found in [3]. Relevant background for this problem in the dynamic setting, including background on its worst-case lower bounds and the bounded-degree case, can be found in [24].
In the Bipartite Perfect Matching problem, the graph is bipartite, and both of its parts have the same size $n$. The goal is to decide whether the maximum matching has cardinality $n$ or not. A recent algorithm for this problem solves it in near-linear time [14], but it remains hard in the dynamic setting. Conditional lower bounds and background for this problem can be found in [4] and [24].
$k$-Clique Detection, $k$-Clique Counting, and Max-Clique.
In the $k$-Clique Detection problem, the goal is to decide whether a graph contains a $k$-clique as a subgraph. In the $k$-Clique Counting problem, the goal is to count the number of copies of a $k$-cliques in a graph. Both of these problems are fundamental “hard” problems in fine-grained complexity. See for reference [43] and [1]. In the Max-Clique problem, the goal is to compute that largest $k$ such that $G$ contains a $k$-clique as a subgraph. This is one of the most well-known NP-hard problems. See [46] for the exponential complexity of this problem (presented for the equivalent MIS problems). I
Densest Subgraph.
In the Densest Subgraph problem, we define the density of a subset $\emptyset\neq S\subseteq V$ as $\rho(S):=e(S)/|S|$, where $e(S)$ is the number of edges in the subgraph induced by $S$. The goal is to find the maximum density of any subset in the graph. The goal is to compute the maximum density of any induced subgraph in a graph. In the static setting, this problem can be solved via $O(\log n)$ max-flow computations,  [21], which results in a near-linear time algorithm [14]. However, it remains hard in the dynamic setting. See [24] for some background and a lower bound. In the Densest Subgraph ($m\geq\alpha n$) problem, we restrict the input graph to have at least $\alpha n$ edges.
$H$-Subgraph Detection for $H$ without pendant vertices.
See [3] for the definition and background for this problem.
Max-Cut.
In this problem, the goal is to compute $\max_{S\subseteq V}e(S,V\setminus S)$. This problem is one of the most well-known NP-hard problems. The fastest exponential-time algorithm for it is by Williams [44].
$st$-SP.
Given a dynamic graph with two distinguished vertices $s$ and $t$, maintain the length of a shortest path between them in the graph. See [4] and [24] for background on this problem.
Online Matrix-vector Multiplication Problem (OMv).
[[24, 45]] In this problem, an algorithm is given an $n$-by-$n$ boolean matrix $M$, preprocesses is in polynomial time, and then it receives a sequence of $n$ vectors $v_{1},v_{2},\ldots,v_{n}\in\{0,1\}^{n}$ one by one. The goal is to compute the boolean product $Mv_{i}$ for every $i$ before seeing the next vector in the sequence. See [24] for background on this problem, where they also propose the OMv Conjecture:
Conjecture 1 (OMv Conjecture [24]).
For any constant $\varepsilon>0$, there is no $O(n^{3-\varepsilon})$-time randomized algorithm that solves OMv with an error probability of at most $1/3$.
In addition, the authors also prove that the OMv problem is essentially equivalent to the OuMv problem, where instead of a sequence of vectors, the algorithm receives a sequence of pairs of vectors $(u_{1},v_{1}),\ldots,(u_{n},v_{n})$, and the goal is to compute the boolean product $u_{i}^{T}Mv_{i}$ for every $i$.